### v5

# Privacy-Aware Sequential Learning

[…]
Privacy-Aware Sequential Learning
[…]
Privacy-Aware Sequential Learning
[…]
In settings like vaccination registries, individuals act after observing others, and the resulting public records can expose private information. We study privacy-preserving sequential learning, where agents add endogenous noise to their reported actions to conceal private signals. Efficient social learning relies on information flow, seemingly in conflict with privacy. Surprisingly, with continuous signals and a fixed privacy budget $(\varepsilon)$, the optimal randomization strategy balances privacy and accuracy, accelerating learning to $\Theta_{\varepsilon}(\log n)$, faster than the nonprivate $\Theta(\sqrt{\log n})$ rate. In the nonprivate baseline, the expected time to the first correct action and the number of incorrect actions diverge; under privacy with sufficiently small $\varepsilon$, both are finite. Privacy helps because, under the false state, agents more often receive signals contradicting the majority; randomization then asymmetrically amplifies the log-likelihood ratio, enhancing aggregation. In heterogeneous populations, an order-optimal $\Theta(\sqrt{n})$ rate is achievable when a subset of agents have low privacy budgets. With binary signals, however, privacy reduces informativeness and impairs learning relative to the nonprivate baseline, though the dependence on $\varepsilon$ is nonmonotone. Our results show how privacy reshapes information dynamics and inform the design of platforms and policies.
[…]
sequential learning, privacy-preserving learning, information cascade, learning efficiency
1 Introduction
When making decisions, people often combine their private information with public information, such as observable behaviors of others (Bikhchandani et al. 2024). For example, to decide to receive a vaccine, people can consult the scientific evidence on the efficacy and safety of the vaccine and weigh this information against the decisions and experiences of others in their community. This process can be modeled within the framework of sequential learning to understand how individual decisions evolve over time based on the actions of their predecessors (Banerjee 1992, Bikhchandani et al. 1992, Acemoglu et al. 2011).
Traditionally, sequential learning models examine how individuals infer from others’ actions, typically assuming that these actions are not distorted by endogenous noise (Ali 2018). While such models have been successful in explaining herding behaviors and information cascades, they overlook privacy concerns that are particularly salient in sensitive domains (e.g., health or politics). In these settings, individuals may refrain from participating, mute unpopular views, or avoid choices that could reveal personal experiences. In the vaccine adoption example, people interpret clinical evidence and personal anecdotes through the lens of their medical histories (e.g., autoimmune conditions), making private signals highly sensitive. Although such information is not directly observable, the sequential nature of adoption—where individuals act after observing others—creates opportunities for inference. This is especially true in public or semi-public contexts (e.g., workplace or school immunization records), where actions can inadvertently reveal private health details. Anticipating this, individuals may strategically misreport their actions to protect privacy, biasing the public record, confounding perceived confidence, and ultimately undermining collective decision quality.
However, randomization can sometimes be beneficial: occasional disregard for predecessors’ actions can prevent fragile, self-reinforcing cascades (Peres et al. 2020); similarly, privacy-motivated noise can dilute early trends driven by misinformation, social pressure, or institutional incentives, reducing misleading cascades and promoting more robust learning. This interplay between privacy and learning efficiency raises a central question: What are the implications of endogenous privacy-preserving behavior for sequential learning performance?
To address these questions, we adopt the metric differential privacy (mDP) framework (Chatzikokolakis et al. 2013) to study the behavior of privacy-aware agents. In our setting, agents endogenously add noise to their actions to limit inferences from other observers, thereby altering the dynamics of learning. Contrary to the common intuition that additional noise should reduce the informativeness of public observations (Le et al. 2017, Papachristou and Rahimian 2025), we show that privacy-motivated randomization can enhance learning by asymmetrically amplifying the log-likelihood ratio across states, thus increasing the information conveyed by each action. This asymmetry stems from a strategic tradeoff between privacy and utility: agents randomize more when their actions would otherwise reveal sensitive information, and less when the risk of revelation is low. Unlike prior work that ties learning speed solely to the structure of private signals (Hann-Caruthers et al. 2018, Rosenberg and Vieille 2019), we demonstrate that privacy-driven noise reshapes aggregation dynamics and can accelerate convergence.
These findings carry important managerial implications for platforms and policymakers. By identifying the mechanisms through which privacy-motivated randomization accelerates learning, designers can tailor data-collection protocols to the sensitivity of the information and the stage of the learning process—for example, adjusting the frequency and granularity of data requests when privacy risks are high, or selectively soliciting input from less-represented groups to counteract early cascades. In data aggregation, incorporating models that account for the presence and strategic use of noise can help platforms weight incoming signals appropriately, filter out spurious early trends, and reinforce robust patterns. Such adaptive strategies can be applied in domains such as consumer-behavior analytics, public-health surveillance, and policy consultation, enabling faster and more reliable collective decision-making while safeguarding individual privacy.
1.1 Main related work
Our work contributes to the literature on sequential learning, particularly regarding whether asymptotic learning occurs and how fast it proceeds. The failure of asymptotic learning due to information cascades was first introduced by Banerjee (1992) and Bikhchandani et al. (1992), who showed that binary private signals can trigger cascades that stop further learning. Smith and Sørensen (2000) later established that unbounded private beliefs are necessary for asymptotic learning, a result extended by Acemoglu et al. (2011) to network settings with partial observations. Subsequent studies confirmed that learning persists even when agents observe random predecessors without time indices (Smith and Sorensen 2013). Where earlier studies consider environments without endogenous distortions, we examine how strategically introduced privacy noise reshapes these learning outcomes. In the binary-signal setting, we find that the probability of a correct cascade is non-monotonic in the strength of privacy concerns: stronger privacy can, in some cases, improve performance by helping the public escape entrenched cascades. When signals are continuous, privacy-aware agents follow a smooth randomized response strategy that balances accuracy with privacy protection, thereby sustaining asymptotic learning—and, counterintuitively, doing so at a faster asymptotic rate then the non-private baseline described below.
Beyond the question of whether asymptotic learning occurs, another strand of research studies its speed. Chamley (2004, Chapter 4) investigates convergence rates in Gaussian settings through computational simulations. Rosenberg and Vieille (2019) analyze the time until the first correct action appears and the total number of incorrect actions, deriving conditions under which these quantities have finite expectations. Hann-Caruthers et al. (2018) precisely characterize the asymptotic speed of learning, showing that for Gaussian signals, the log-likelihood ratio evolves at a rate of $\Theta(\sqrt{\log n})$—so slow that the expected time for learning to begin is infinite, even though asymptotic learning is guaranteed. Arieli et al. (2025) examine the role of overconfidence, demonstrating that mildly overconfident agents, who underestimate the quality of their peers’ information, can accelerate learning in standard sequential models. We extend this literature by incorporating privacy-aware agents and showing how endogenous noise fundamentally reshapes both the dynamics and the speed of learning. Under a homogeneous privacy budget regime, rational agents adopt smooth randomized response strategies that balance privacy and utility, leading to faster asymptotic learning at rate $\Theta_{\varepsilon}(\log n)$—a sharp improvement over the $\Theta(\sqrt{\log n})$ rate in Hann-Caruthers et al. (2018). In heterogeneous populations, the presence of agents with very small privacy budgets (approaching zero) further accelerates information aggregation, yielding an order-optimal asymptotic learning rate of $\Theta(\sqrt{n})$. Moreover, under sufficiently strong privacy concerns, both the expected time to the first correct action and the expected number of incorrect actions remain finite. These gains stem from the asymmetric amplification of correct signals induced by strategic privacy noise, revealing that privacy-aware behavior can enhance—rather than hinder—learning efficiency.
Our study contributes to the broader literature on how privacy constraints affect learning and decision-making. In the context of sequential learning, Tsitsiklis et al. (2021) and Xu (2018) analyze the query complexity of private learning and show that stronger privacy constraints require more information acquisition for accurate inference. Although their models share structural similarities with ours, they focus on independent queries, whereas we study path-dependent decision-making in which each agent’s action shapes the information available to others. In distributed settings, privacy-preserving noise is typically modeled as exogenous (Papachristou and Rahimian 2025, Rizk et al. 2023, Tao et al. 2023), whereas we treat it as an endogenous choice. We further allow agents to differ in their privacy budgets, consistent with evidence of heterogeneous privacy preferences (Acquisti and Grossklags 2005) and with personalized/heterogeneous DP frameworks (Alaggan et al. 2015, Acharya et al. 2025). This shift from algorithmic to behavioral modeling highlights how rational agents strategically balance privacy and learning performance, offering new insights at the intersection of privacy theory and social learning. To our knowledge, the only work examining social learning in privacy-preserving data collection is Akbay et al. (2021), which focuses on simultaneous actions and does not address sequential learning or information cascades.
Our work is the first to examine how privacy concerns shape sequential learning with endogenous noise. Counter to the common intuition drawn from exogenous-noise models, we find that endogenous privacy noise does not necessarily harm learning performance. Most privacy-preserving designs ignore the dynamic nature of decision-making and inject fixed, excessive noise based on global sensitivity bounds. In contrast, endogenous noise adapts to the evolving decision process, adjusting in response to both privacy concerns and utility considerations, thereby improving learning efficiency. Further related literature is reviewed in Appendix 6.
1.2 Main contribution
In this paper, we introduce a structured approach to analyzing privacy-aware sequential learning, offering new theoretical insights into learning efficiency under privacy concerns. When private signals are binary, we reformulate the problem as a Markov chain with absorbing states and establish results on the probability of a correct cascade, showing how privacy concerns alter the likelihood of learning from past observations. Using techniques from the gambler’s ruin problem, we quantify the effect of privacy-related noise on belief propagation and show that such noise can systematically reduce the probability of correct cascades.
In the continuous signal case, rational agents adopt a smooth randomized response strategy that selectively limits the amount of noise based on the potential for privacy loss. This strategy allows agents to protect their privacy while minimizing unnecessary randomization that would otherwise impair learning. We then analyze the evolution of the log-likelihood ratio over time using Gaussian tail approximations and differential equations, which enables a rigorous characterization of how information accumulates under endogenous privacy behavior. Actions shaped by privacy concerns more accurately reflect a combination of private signals and observed history, leading to an accelerated asymptotic learning speed of $\Theta_{\varepsilon}(\log(n))$, in contrast to the classical $\Theta(\sqrt{\log(n)})$ rate in the non-private regime (Rosenberg and Vieille 2019, Hann-Caruthers et al. 2018). The increase in asymptotic learning speed arises from the inherent asymmetry of the smooth randomized response strategy under different states. Specifically, under the true state, the average probability of flipping an action is lower than under the false state. As a result, agents’ actions become more informative during the inference process, enabling the public belief to update more efficiently. To better evaluate the efficiency of sequential learning under privacy constraints, we also consider two additional measures: the expected time to the first correct action and the expected total number of incorrect actions. We find that both quantities remain finite when privacy concerns are strong, which stands in contrast to traditional results in the non-private setting, where both expectations are known to diverge (Rosenberg and Vieille 2019, Hann-Caruthers et al. 2018). Furthermore, based on numerical simulations, we identify the existence of an optimal privacy budget $\varepsilon^{*}$ that minimizes both the expected time to the first correct action and the expected number of incorrect actions. This highlights a fundamental trade-off between privacy and learning efficiency.
Learning Metrics Non-private Private-Homogeneous Private-Heterogeneous
Binary
Probability of correct cascades $\frac{\rho^{k}-1}{\rho^{2k}-1}$ $\frac{\rho(\varepsilon)^{k}-1}{\rho(\varepsilon)^{2k}-1}$ $\frac{\bar{\rho}^{k}-1}{\bar{\rho}^{2k}-1}$
Information cascade threshold $k=2$ $k=\left\lfloor\log_{\rho(\varepsilon)}\frac{1-p}{p}\right\rfloor+1$ $k=\left\lfloor\log_{\bar{\rho}}\frac{1-p}{p}\right\rfloor+1$
Continuous
Convergence rate $\Theta(\sqrt{\log n})$ $\Theta_{\varepsilon}(\log n)$ $\Theta(\sqrt{n})$
Time to first correct action $C_{1}\sum_{n=1}^{\infty}e^{-\frac{2\sqrt{2}}{\sigma}\sqrt{\log n}}$ $C_{1}C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}e^{-\frac{2}{\varepsilon\sigma^{2}}\log{n}}$ $C_{1}\sum_{n=1}^{\infty}e^{-\tilde{C}n^{1/2}}$
Table 1: Summary of theoretical results under different privacy settings. Here, $\rho=\frac{1-p}{p},\rho(\varepsilon)=\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}$, $\bar{\rho}=\frac{(1-\bar{u})(1-p)+\bar{u}p}{\bar{u}(1-p)+p(1-\bar{u})}$, $\bar{u}=E_{\varepsilon}[u(\varepsilon)]=E_{\varepsilon}\left[\frac{1}{1+e^{\varepsilon}}\right]$, $C(\varepsilon)=\frac{\varepsilon\sigma^{2}}{4}\left(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}\right)$, and $C_{1}$ and $\tilde{C}$ are constants. The total number of incorrect actions differs from the time of the first correct action by a constant factor and is omitted.
We also extend our analysis to the heterogeneous setting by considering the case where the privacy budget $\varepsilon$ follows a uniform distribution $\varepsilon\sim U(0,1)$. Due to the presence of agents with strong privacy concerns (i.e., small $\varepsilon$), we show that the convergence rate of the public log-likelihood ratio can increase to $\Theta(\sqrt{n})$. Furthermore, we establish that this is the optimal convergence rate that can be achieved under any distribution of privacy budgets. This result offers two key insights. First, it demonstrates that the structured randomness introduced by privacy concerns can actually accelerate the learning process. In contrast to the non-private sequential learning setting—where the asymptotic convergence rate is only $\Theta(\sqrt{\log(n)})$ (Rosenberg and Vieille 2019, Hann-Caruthers et al. 2018)—the heterogeneous private setting achieves a significantly faster rate. Second, although privacy-aware sequential learning benefits from increased speed, it still falls short of the ideal scenario in which agents directly observe i.i.d. signals. In the latter case, the convergence rate reaches the optimal order of $\Theta(n)$, as guaranteed by the law of large numbers. Thus, while privacy can be compatible with efficient learning, sequential learning inherently limits the maximal speed of information aggregation compared to fully transparent settings. A comprehensive comparison of these results across different privacy models is summarized in Table 1.
Our analysis demonstrates that, contrary to conventional belief, privacy-aware behavior can enhance learning by altering the dynamics of information flow over time. Unlike prior literature, which primarily focuses on asymptotic learning without privacy considerations, we provide the first formal convergence rate analysis under endogenous privacy concerns. These findings reveal how strategic privacy behavior interacts with sequential learning to shape collective decision outcomes. From a managerial perspective, this highlights that respecting individual privacy preferences need not come at the cost of slower learning—well-structured privacy practices can actually accelerate collective insight and support more informed decision-making over time.
The remainder of this paper is organized as follows. First, we formally define our problem setup in Section 2, including the sequential decision making framework and integration of privacy constraints. Section 3 focuses on the case of binary signals, where we analyze how differential privacy affects the probability of correct cascades using a Markov chain representation. In Section 4, we extend our analysis to continuous signals, introducing the smooth randomized response strategy and showing how it accelerates and improves learning efficiency. Finally, Section 5 summarizes our findings and discusses potential future research directions, including broader implications for privacy-aware decision making in sequential learning environments.
2 A Model of Sequential Learning with Privacy
We consider a classical sequential learning model: There exists an unknown binary state of nature, denoted by $\theta\in\{-1,+1\}$, with the same prior probability $1/2$; each agent, indexed by $n\in\{1,2,\ldots\}$, receives a private signal $s_{n}\in\mathscr{X}$, drawn independently and identically distributed (i.i.d.) from a conditional distribution $F_{\theta}$, that is, $s_{n}\sim F_{\theta}$, where $F_{\theta}$ depends on the true state $\theta$. Without considering privacy concerns, each agent $n$ takes an action $a_{n}\in\{-1,+1\}$ based on the information received, including the history of previous reported actions denoted by $h_{n-1}:=\{x_{1},\cdots,x_{n-1}\}$, where $x_{i}$ is the reported action of agent $i$, and their own private signal $s_{n}$.
With the introduction of privacy concerns, agents experience privacy loss when reporting their actions. In the context of vaccine adoption, the binary state $\theta$ represents whether the vaccine is truly effective and safe for the general population, where $\theta=+1$ indicates effectiveness and $\theta=-1$ indicates ineffectiveness or harm. Each agent $n$ receives a private signal $s_{n}$, which reflects their individualized assessment based on clinical research, anecdotal reports, and personal health considerations. For instance, someone with a preexisting autoimmune condition may investigate whether the vaccine triggers flare-ups or remains effective under such conditions. These signals are thus not only informative but also intimately tied to an individual’s sensitive medical history. Publicly revealing one’s vaccine decision may inadvertently disclose these private health concerns, leading to potential privacy loss. As a result, agents must balance the value of contributing to collective knowledge with the personal cost of revealing information about their health.
To systematically characterize the extent of privacy concerns in a sequential learning process, we introduce the notion of a privacy budget. Each agent is associated with a privacy budget that reflects their level of privacy concern: agents with stronger privacy concerns have smaller privacy budgets, resulting in noisier reported actions; in contrast, agents with weaker privacy concerns have larger privacy budgets and are more likely to report their actions truthfully. When making decisions, agents ensure that their chosen actions do not violate their individual privacy budgets.
To formalize this idea, we adopt the metric differential privacy framework (Chatzikokolakis et al. 2013), which allows us to rigorously quantify privacy guarantees in settings with continuous inputs. The privacy budget $\varepsilon$ quantifies how distinguishable an agent’s private signal is under the randomized reporting strategy $\mathcal{M}$, and serves as a key parameter in understanding the trade-off between privacy and information aggregation. For simplicity, we assume a homogeneous privacy budget across agents in the main analysis. A discussion of the heterogeneous case is provided in Section 3 and  4 .
Definition 2.1 ($(\varepsilon,d_{\mathscr{X}})$-mDP for the Sequential Learning Model)
Let $\mathscr{X}$ be a set equipped with a metric $d_{\mathscr{X}}:\mathscr{X}\times\mathscr{X}\to[0,\infty)$. A randomized strategy $\mathcal{M}$ satisfies $(\varepsilon,d_{\mathscr{X}})$-metric differential privacy if for all $s_{n},s_{n}^{\prime}\in\mathscr{X}$ and all possible outputs $x_{n}\in\{-1,+1\}$, we have:
$\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})\leq\exp\left(\varepsilon\cdot d_{\mathscr{X}}(s_{n},s_{n}^{\prime})\right)\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n}).$ (1)
Here, $\varepsilon>0$ is the privacy budget.
In the context of vaccine adoption, the choice of $\varepsilon$ determines the level of noise added to the reported actions. A high privacy budget implies limited randomization, making the observed adoption statistics closely reflect individuals’ true decisions, and thus a reliable indicator of public confidence in the vaccine. In contrast, a lower privacy budget introduces greater randomness, which obscures individual-level choices and helps protect sensitive health information. This reduces the risks of social pressure, stigmatization, or strategic behavior related to personal medical conditions, thereby improving privacy at the potential cost of reduced inference accuracy.
This privacy-driven distortion affects the information aggregation process in sequential vaccine uptake. If early adopters misreport their actions, later individuals—who rely on observed behavior to update their beliefs—may be misled, compounding initial biases and undermining accurate social learning. However, privacy-preserving noise can also introduce beneficial uncertainty. By disrupting deterministic adoption patterns, it can reduce the risk of irrational herding, especially in situations where early adoption is influenced by misinformation, social pressure, or unrepresentative signals. In such cases, randomization can prevent premature consensus and preserve the diversity of private information, ultimately improving long-run aggregation accuracy. This highlights the dual role of privacy in sequential learning: Although it can distort belief formation and bias perception of public consensus, it can also protect against overconformity and foster more resilient and balanced collective decision making.
Based on their private signal and the observed history of prior adoption decisions, each agent chooses a true adoption decision $a_{n}$. To protect privacy, the agent may then misreport via a randomized mechanism $\mathcal{M}$ that injects bounded noise into the publicly observed action. This reporting is not fully random: agents internalize the value of informative records—for example, for eligibility verification, booster scheduling, adverse-event monitoring, and care coordination—and they may face moral, reputational, or regulatory costs from blatant misreporting. Thus, If the reported action $x_{n}$ matches $\theta$, the agent’s utility is 1 ; otherwise, it is 0. Accordingly, $\mathcal{M}$ is calibrated by a fixed privacy budget to balance privacy concerns with public informativeness. Formally, agent $n$ selects a reporting strategy to maximize their expected utility, subject to a fixed privacy budget:
$\mathcal{M}_{n}^{*}=\arg\max_{\mathcal{M}\in\mathcal{M}_{\varepsilon,d_{\mathscr{X}}}}\,\mathbb{E}\left[u_{n}(x_{n},\theta)\mid h_{n-1},s_{n}\right],$
where $x_{n}\sim\mathcal{M}(a_{n})$ denotes the reported (randomized) action, $a_{n}$ is the true intended action of agent $n$, $h_{n-1}$ is the public history observed up to time n1n-1, and $\mathcal{M}_{\varepsilon,d_{\mathscr{X}}}$ denotes the set of strategies that satisfy $(\varepsilon,d_{\mathscr{X}})$-metric differential privacy.
3 Binary Signals and Information Cascades
In the binary model, we assume that $s_{n}\in\mathscr{X}$, where $\mathscr{X}=\{-1,+1\}$, and that $s_{n}=\theta$ with probability $p>1/2$. That is, each private signal matches the true state $\theta\in\{-1,+1\}$ with probability $p$, and equals $\theta$ with probability 1p1-p. In this binary signal setting, we endow the signal space $\mathscr{X}$ with the Hamming distance $d_{h}$, which takes the value 0 if two signals are equal and 1 otherwise. Under this choice of metric, metric differential privacy coincides with local differential privacy Chatzikokolakis et al. (2013). The formal definition is as follows:
Definition 3.1 ($\varepsilon$-Local Differential Privacy for the Binary Model)
A randomized strategy $\mathcal{M}$ satisfies $\varepsilon$-local differential privacy if for all $s_{n},s_{n}^{\prime}\in\{-1,+1\}$ and for all possible outputs $x_{n}\in\{-1,+1\}$, we have:
$\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})\leq\exp(\varepsilon)\cdot\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n}).$ (2)
Here, $\varepsilon>0$ is the privacy budget.
In this paper, we show that the optimal reporting strategy under a fixed privacy budget takes the form of a randomized response. This means that each agent may randomly alter their action to conceal their private signal. For example, given the observational history and private signal, an agent might ideally choose action $a_{n}=+1$. However, due to a limited privacy budget, they may instead report 1-1 with probability $u_{n}$. We denote the action of agent $n$ after applying the randomized response by $x_{n}$. The differential privacy requirement for the randomized response strategy in the binary setting is as follows.
Definition 3.2 (Randomized Response Strategy)
A strategy $\mathcal{M}(s_{n};h_{n-1})$ is called a randomized response strategy with flip probability $u_{n}$ if, given an initial decision $a_{n}$, the reported action $x_{n}$ is determined as follows:
$\mathbb{P}_{\mathcal{M}}(x_{n}=+1\mid a_{n}=-1)=\mathbb{P}_{\mathcal{M}}(x_{n}=-1\mid a_{n}=+1)=u_{n},$
$\mathbb{P}_{\mathcal{M}}(x_{n}=+1\mid a_{n}=+1)=\mathbb{P}_{\mathcal{M}}(x_{n}=-1\mid a_{n}=-1)=1-u_{n}.$
Here, $a_{n}$ represents the initial action of the agent, which is based on their private signal $s_{n}$ and the observed history $h_{n-1}$. The randomized response strategy then perturbs this action according to probability $u_{n}$, flipping it with probability $u_{n}$ and keeping it unchanged with probability $1-u_{n}$.
3.1 Main Results for Binary Signals
First, we present a significant finding that greatly simplifies the analysis of sequential learning with privacy constraints in the binary case. Before proceeding, we define an information cascade, which is important for simplification.
Definition 3.3 (Information Cascade)
An information cascade occurs when all agents herd on the public belief, meaning that each agent’s action becomes independent of their private signal. If actions taken after the onset of the information cascade are aligned with the true state, it is referred to as a correct cascade. In contrast, if actions are misaligned with the true state, it is called a wrong cascade.
The following theorem shows that, prior to the initiation of the information cascade, different agents use the same randomized response strategy with constant probability $u_{n}=u(\varepsilon)=\frac{1}{1+e^{\varepsilon}}$. Once the information cascade occurs, the agents do not need to add noise to their actions, so $u_{n}=0$.
Theorem 3.4 (Randomized Response Strategy for the Binary Model)
Before the occurrence of an information cascade, the optimal reporting strategy for each agent takes the form of a randomized response, with probability $u_{n}=u(\varepsilon)=\frac{1}{1+e^{\varepsilon}}$. Once an information cascade occurs, agents report their actions truthfully.
Before an information cascade occurs, each agent’s action is solely determined by their private signal, meaning that their decisions directly reflect the information they receive. Specifically, if an agent receives a signal $s_{n}=+1$, they will choose action $a_{n}=+1$ before applying the randomized response strategy. To protect their private signals from being inferred by others, agents must use the same randomized response strategy, introducing controlled noise into their actions while maintaining the statistical properties of the decision process.
To illustrate, consider the vaccine adoption scenario discussed earlier, in which individuals decide sequentially whether to take a new vaccine, and each person observes aggregate adoption statistics before making their decision. In the early stages, individuals rely on their private health-related signals and can apply randomized response techniques to protect sensitive medical information, such as reporting adoption with probability $u_{n}$ even if they privately choose not to adopt and vice versa. This strategy ensures plausible deniability and guards against inference of personal health conditions, especially when the vaccine’s risks or effectiveness are controversial.
However, once an information cascade sets in, the dynamics shift fundamentally. Individuals no longer rely on their private signals to decide, but instead follow the observed behavior of earlier adopters. In this context, if early adoption statistics strongly indicate widespread adoption, later individuals can conform to the trend, assuming that it reflects collective medical consensus, regardless of their personal concerns.
At this point, actions are no longer informative about private signals, and thus adding noise for privacy protection becomes unnecessary. This marks a key transition in privacy-preserving sequential learning: before the cascade, privacy-aware strategies are crucial to protecting sensitive personal information; after the cascade, such strategies may be redundant, as decisions are fully driven by public signals.
While our main analysis focuses on agents who endogenously choose privacy-preserving noise, our results also have implications for systems in which noise is applied exogenously. In particular, public health platforms that release vaccine uptake data in real time may implement privacy strategies to protect sensitive individual information. If such noise continues to be applied after a cascade has already formed, it may introduce unnecessary distortion without improving privacy. Adaptive privacy strategies that adjust based on whether a cascade has emerged could better balance the trade-off between protecting individual privacy and preserving the integrity of social learning. This underscores the importance of dynamic privacy strategies in sequential public health decision making, especially when vaccine adoption is shaped by social sensitivity or misinformation.
To further analyze the impact of the differential privacy budget on learning performance, we need to define the threshold for the information cascade as follows:
Definition 3.5 (Information Cascade Threshold $(k)$)
The information cascade threshold $k$ is defined as the difference between the occurrences of action +1+1 and action -$1$ in the observation history $h_{n-1}$ that makes agent $n$ indifferent to their private signals for any $n\geq k$. If the difference between the occurrences of the action +1+1 and the action 1\-1 in $h_{n-1}$ exceeds $k$, then agent $n$ will choose the action +1+1 regardless of their private signal. In contrast, if the difference between the occurrences of the action 1-1 and the action +1+1 in $h_{n-1}$ exceeds $k$, then agent $n$ will choose the action 1-1 regardless of their private signal.
Consider the case where no privacy-preserving noise is introduced, i.e., $\varepsilon=\infty$. Without noise, it was shown in Bikhchandani et al. (1992) that cascading begins with the first agent $n$ for whom the observed history satisfies $k=2$, meaning that the agent observes one action at least twice more frequently than the other. That is, if an agent sees two more occurrences of action +1+1 than 1-1 in $h_{n-1}$, they will always choose +1+1, even if their private signal suggests otherwise. Similarly, if an agent observes two more occurrences of action 1-1 than +1+1, they will always choose 1-1, leading to an irreversible information cascade.
Theorem 3.6 (Probability of the Correct Cascade)
Consider a correct signal probability $p$ and privacy budget $\epsilon$, and define $\alpha=(1-p)/p$ and $u(\varepsilon)=\frac{1}{1+e^{\varepsilon}}$. Let $v_{k}=\frac{1-\alpha^{\frac{k-2}{k-1}}}{1-\alpha^{\frac{k-2}{k-1}}+\alpha^{\frac{-1}{k-1}}-\alpha}$ and $\varepsilon_{k}=\log\frac{1-v_{k}}{v_{k}}$, where $k$ is the information cascade threshold given by $k=\lfloor\log_{\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}}\frac{1-p}{p}\rfloor+1$. Subject to the privacy budget $\varepsilon$, agents employ the randomized response strategy with flip probability $u_{n}=u(\varepsilon)$; subsequently, for $\varepsilon\in(\varepsilon_{k+1},\varepsilon_{k}]$, the information cascade threshold $k$ does not change and the probability of the correct cascade increases with $\varepsilon$.
In the non-private sequential learning model with binary signals, the information cascade threshold $k$ is only a function of the probability $p$ of receiving correct signals. However, the randomized response strategy affects the information cascade threshold $k$, which can vary with the privacy budget $\varepsilon$ for a fixed $p$. Without using randomized response, as long as an information cascade is not started, an agent receiving a signal of $s_{n}=+1$ will choose the action $a_{n}=+1$ with probability one and the action $a_{n}=-1$ with probability zero. However, following the randomized response strategy, the probability difference between the choice of different actions narrows, reducing the informational value that each randomized action $x_{n}$ conveys. Consequently, a larger difference in the number of actions is required to trigger an information cascade. Given an information cascade threshold $k$, for $\varepsilon\in(\varepsilon_{k+1},\varepsilon_{k}]$, the information cascade threshold $(k)$ does not change with increasing privacy budget ($\varepsilon$). In this range, $\varepsilon\in(\varepsilon_{k+1},\varepsilon_{k}]$, as the privacy budget increases, the flip probability $u$ of the randomized response strategy decreases, the agents are more inclined to select the correct actions and their actions contain more information about their private signals. Consequently, the probability of the correct cascade will increase. However, increasing $\varepsilon$ beyond $\varepsilon_{k}$ will cause the information cascade threshold to decrease by $1$ from $k$ to k1k-1. With the information cascade occurring earlier for $\varepsilon\in(\varepsilon_{k},\varepsilon_{k-1}]$ there is less opportunity for information aggregation and we see a drop in the probability of correct cascades by increasing $\varepsilon$ above $\varepsilon_{k}$ at the end of each $(\varepsilon_{k+1},\varepsilon_{k}]$ interval; see Figure 1. It is worth mentioning that this nonmonotonic pattern of cascade probability as a function of $\varepsilon$ also has important implications for settings where privacy budgets are externally chosen. Although our main model assumes that agents choose noise levels endogenously, this result suggests that when a platform or policymaker selects a fixed privacy budget for a population, the endpoints $\varepsilon_{k}$ of the stability intervals are normatively attractive. Specifically, each $\varepsilon_{k}$ dominates nearby higher values of $\varepsilon$ in terms of privacy protection and learning performance. This offers a principled guide for setting privacy budgets in binary decision-making environments.
Figure 1: Probability of correct cascade vs. privacy budget ($\varepsilon$) for different cascade thresholds ($k$). Each colored line represents the probability of a correct cascade for a specific threshold $k$.
Figure 1 illustrates how the probability of a correct cascade varies with the differential privacy budget $\varepsilon$ across different cascade thresholds $k$, considering three values of $p$: $p=0.55$, $p=0.7$, and $p=0.9$. Each colored line represents a specific value of $k$, with vertical dashed lines indicating the transition points between thresholds. In particular, the probability of a correct cascade does not increase monotonically with $\varepsilon$; at certain points, an increase in $\varepsilon$ leads to a reduction in the cascade threshold, which in turn decreases the probability of a correct cascade. This effect is more pronounced at higher values of $p$; For example, when $p=0.9$, the drop in the probability of a correct cascade during the transition from $k=2$ to $k=3$ is more substantial than that for $p=0.55$.
To better understand these non-monotonic behaviors and their limiting cases, we turn to the analytical expressions that govern the cascade threshold and the corresponding probability of a correct cascade. The probability of a correct cascade is derived in Lemma 8.2, where it is shown that
$\mathbb{P}(\text{correct cascade})=\frac{\rho(\varepsilon)^{k}-1}{\rho(\varepsilon)^{2k}-1},$
where $\rho(\varepsilon)=\frac{(1-u(\varepsilon))(1-p)+pu(\varepsilon)}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}$. The associated cascade threshold is given in Lemma 8.1:
$k=\left\lfloor\log_{\rho(\varepsilon)}\frac{1-p}{p}\right\rfloor+1.$ (3)
At the boundary of each interval—that is, as $\varepsilon\to\varepsilon^{+}_{k}$—the threshold becomes asymptotically exact, and the floor operation can be dropped. In this limiting case, we have
$\lim_{\varepsilon\to\varepsilon^{+}_{k}}k=\log_{\rho(\varepsilon^{+}_{k})}\frac{1-p}{p},\quad\lim_{\varepsilon\to\varepsilon^{+}_{k}}\mathbb{P}(\text{correct cascade})=p$
This confirms that the probability of a correct cascade, as $\varepsilon$ approaches the threshold from the right, converges to the signal accuracy.
The Binary Signals Model with Heterogeneous Privacy Budgets
Individuals have been shown to exhibit highly diverse privacy attitudes and behaviors (Acquisti and Grossklags 2005, Jensen et al. 2005). This heterogeneity motivates moving beyond a uniform, exogenous treatment of privacy toward frameworks that allow differentiated, endogenous protection across users. The notion of heterogeneous differential privacy (Alaggan et al. 2015) and personalized variants such as individualized budgets for regression tasks (Acharya et al. 2025) embody this perspective by assigning each individual their own privacy parameter rather than enforcing a single level across the population. Such heterogeneity is particularly relevant in contexts where privacy valuations directly shape participation decisions. For example, in the case of vaccination data, some individuals may be willing to disclose sensitive information with relatively weak guarantees if doing so accelerates collective immunity, while others may insist on much stronger safeguards before consenting to share their records.
Formally, We extend our model to incorporate heterogeneous privacy budgets. In the binary signal setting, we assume that each agent’s privacy budget $\varepsilon_{n}$ is drawn from a distribution such that the expected flipping probability $\bar{u}=\mathbb{E}_{\varepsilon_{n}}[u_{n}]$ exists, where $u_{n}=\frac{1}{1+e^{\varepsilon_{n}}}$ denotes the probability with which agent $n$ flips their action under the randomized response strategy. For the binary signals model, the key distinction under heterogeneous privacy concerns lies in how agents interpret the information contained in the public history. Since each agent may have a different privacy budget $\varepsilon_{n}$, they adopt different randomized response strategies, resulting in heterogeneous flipping probabilities $u_{n}=\frac{1}{1+e^{\varepsilon_{n}}}$. As a result, when aggregating information from previous actions, agents must consider the expected flipping probability $\bar{u}=\mathbb{E}_{\varepsilon_{n}}[u_{n}]$ rather than a fixed value.
Specifically, when agent $n$ makes her decision, her information consists of both the public history and her private signal. She therefore bases her choice on a combination of the public belief $l_{n}$ and the private belief $L_{n}$. Assuming that an information cascade has not yet occurred, and that there are $k$ more agents who have chosen action +1+1 than those who have chosen 1-1, the public belief is given by:
$l_{n}=\log\frac{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}\mid\theta=+1)}{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}\mid\theta=-1)}=\log\left[\frac{\bar{u}(1-p)+p(1-\bar{u})}{(1-p)(1-\bar{u})+\bar{u}p}\right]^{k}.$
Now, compared to the homogeneous case, the key difference lies in the fact that $\bar{u}$ represents the expected flipping probability, averaged over the distribution of privacy budgets $\varepsilon_{n}$. This added heterogeneity affects how informative each action is, and consequently influences the dynamics of information aggregation. The following theorem characterizes how the cascade threshold and the probability of a correct cascade depend on $\bar{u}$ in this heterogeneous setting.
Theorem 3.7 (Probability of Correct Cascade under Heterogeneous Privacy Budget)
Consider a correct signal probability $p$ and privacy budget $\varepsilon_{n}$, and define $\alpha=(1-p)/p$ and $\bar{u}=\mathbb{E}_{\varepsilon_{n}}[\frac{1}{1+e^{\varepsilon_{n}}}]$. Let $\tilde{v}_{k}=\frac{1-\alpha^{\frac{k-2}{k-1}}}{1-\alpha^{\frac{k-2}{k-1}}+\alpha^{\frac{-1}{k-1}}-\alpha}$, where $k$ is information cascade threshold given by $k=\left\lfloor\log_{\frac{(1-\bar{u})(1-p)+\bar{u}p}{\bar{u}(1-p)+p(1-\bar{u})}}\frac{1-p}{p}\right\rfloor+1$. Subject to a privacy budget $\varepsilon_{n}$ for each agent, agents adopt the randomized response strategy with flipping probability $u_{n}(\varepsilon_{n})=\frac{1}{1+e^{\varepsilon_{n}}}$, then for $\bar{u}\in[\tilde{v}_{k},\tilde{v}_{k+1})$, the cascade threshold $k$ remains unchanged, and the probability of a correct cascade decreases with expected flipping probability $\bar{u}$.
Similar to the homogeneous case, the probability of a correct cascade in the heterogeneous setting is closely related to the average level of privacy concern in the population. When the cascade threshold $k$ remains fixed, a higher expected flipping probability $\bar{u}$—corresponding to stronger average privacy concerns—leads to a lower probability of a correct cascade. This is because increased randomization makes agents’ actions less informative when the threshold for triggering a cascade does not change. However, as $\bar{u}$ continues to increase and crosses a threshold, the cascade threshold $k$ itself may change, effectively delaying the onset of an information cascade. In this regime, agents rely more on their private signals rather than following the public history, which can help them escape the information trap. As a result, the probability of a correct cascade may increase. This leads to a non-monotonic relationship between the average flipping probability and learning performance.
4 Learning Efficiency with Continuous Signals
Building on the binary model, we now turn to the continuous signal setting to provide a more comprehensive understanding of the sequential observational learning problem under different signal structures. In the continuous model, each agent receives a private signal $s_{n}\in\mathbb{R}$, which is conditionally independent and identically distributed given the true state $\theta\in\{\pm 1\}$, specifically following $s_{n}\sim\mathcal{N}(\theta,\sigma^{2})$. To evaluate learning performance, we first focus on a key criterion commonly used in the literature: asymptotic learning, which refers to agents eventually making the correct decision with probability one (see, e.g., Acemoglu et al. (2011)). We formally define it as follows:
Definition 4.1 (Asymptotic Learning)
We say that asymptotic learning occurs if agents’ actions converge to the true state in probability, i.e., $\lim_{n\to\infty}\mathbb{P}(a_{n}=\theta)=1$.
While standard differential privacy offers strong privacy guarantees by requiring indistinguishability across all adjacent inputs, it can fundamentally obstruct asymptotic learning in continuous models. The key issue lies in how privacy is enforced when the signal space is unbounded: to satisfy $\varepsilon$-DP, the strategy must add a uniform level of noise across all possible input signals, regardless of how different they are. This uniform noise requirement forces agents to randomize their actions—even when their private signals provide highly informative evidence about the true state. As a result, agents cannot fully exploit their private information, and their actions retain a persistent level of randomness throughout the learning process. This prevents the public belief from concentrating around the true state and leads to a failure of asymptotic learning. In other words, the DP constraint injects sufficient uncertainty into every action such that the sequence of decisions does not converge to the correct state with high probability, even as the number of agents grows.
To illustrate the limitations of standard DP in this setting, we show that under an $\varepsilon$-DP constraint, agents are forced to adopt a randomized response strategy with a constant flip probability, regardless of their signal strength or their position in the sequence.
Figure 2: Probability of action +1+1 given signal $s_{n}$ (blue), and signal distribution under $\theta=+1$ (orange).
As illustrated in Figure 2, for each value of the public belief $\ell$, there exists a unique signal threshold $t(\ell)$ such that agent $n$’s optimal action switches from 1-1 to +1+1 as their private signal $s_{n}$ crosses this threshold. This indicates that decisions are most sensitive to noise near this threshold. However, the $\varepsilon$-DP requirement enforces indistinguishability across all signal pairs, which inevitably includes pairs on opposite sides of the decision boundary. As a result, agents must randomize their actions with the same fixed probability, regardless of the signal’s informativeness. We provide the detailed formal analysis in Appendix 9.
This negative result is rooted in the fact that standard DP imposes a uniform level of noise across the entire signal space. In practice, such uniform randomization is overly cautious—agents with signals far from the decision threshold are unlikely to benefit from privacy noise, yet are still required to randomize. To address this, metric differential privacy offers a more flexible and context-aware privacy notion. Rather than treating all signal pairs as equally sensitive, mDP allows the privacy guarantee to degrade with distance. Specifically, noise is added proportional to the similarity between signals, preserving accuracy in robust regions and only perturbing decisions near thresholds. This refinement makes mDP particularly well-suited to continuous settings, where the proximity of signals plays a key role in decision-making. To ensure the robustness of our conclusions, we also examine an alternative privacy notion—pufferfish privacy—which is also suitable for unbounded Gaussian signals. Results under this definition are provided in Appendix 19. We formally define mDP for the continuous setting as follows:
Definition 4.2 ($(\varepsilon,d_{\mathbb{R}})$-mDP for the Continuous Model)
A randomized strategy $\mathcal{M}:\mathbb{R}\to\{-1,+1\}$ satisfies $(\varepsilon,d_{\mathbb{R}})$-metric differential privacy if for all $s_{n},s_{n}^{\prime}\in\mathbb{R}$ and all possible outputs $x_{n}\in\{-1,+1\}$, we have:
$\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})\leq\exp\left(\varepsilon d_{\mathbb{R}}(s_{n},s_{n}^{\prime})\right)\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n}),$ (4)
where $d_{\mathbb{R}}(s_{n},s_{n}^{\prime}):=\|s_{n}-s_{n}^{\prime}\|_{1}$, and $\varepsilon>0$ is the privacy budget.
Given this privacy constraint, agents must carefully design their reporting strategy to balance utility and privacy. In the continuous signal model, agent $n$ selects a randomized strategy from the set $\mathcal{M}_{\varepsilon,d_{\mathbb{R}}}$ that satisfies the $(\varepsilon,d_{\mathbb{R}})$-metric differential privacy condition. The goal is to maximize the agent’s expected utility conditioned on the observed history and the private signal:
$\mathcal{M}_{n}^{*}=\arg\max_{\mathcal{M}\in\mathcal{M}_{\varepsilon,d_{\mathbb{R}}}}\,\mathbb{E}\left[u_{n}(x_{n},\theta)\mid h_{n-1},s_{n}\right],$
where $\mathcal{M}_{\varepsilon,d_{\mathbb{R}}}$ denotes the set of strategies satisfying $(\varepsilon,d_{\mathbb{R}})$-mDP.
While asymptotic learning ensures that agents eventually identify the true state, it does not capture how fast this learning occurs. In practice, the speed at which agents reach correct decisions is often more critical than eventual convergence. To evaluate this aspect, we turn to the notion of learning efficiency, which can be assessed using three commonly adopted measures.
(1) Convergence Rate of Public Beliefs. This measure evaluates how rapidly the public belief converges to the true state as more agents act. The log-likelihood ratio (LLR) of the public belief is defined as:
$l_{n}=\log\frac{\mathbb{P}(\theta=+1\,|\,x_{1},\ldots,x_{n-1})}{\mathbb{P}(\theta=-1\,|\,x_{1},\ldots,x_{n-1})},$ (5)
where $\mathbb{P}(\theta=+1\,|\,x_{1},\ldots,x_{n-1})$ denotes the belief of agent $n$ given the actions of the first n1n-1 agents. We say that $l_{n}$ converges at rate $f(n)$ if the ratio $l_{n}/f(n)$ tends to 1. Formally:
Definition 4.3 (Convergence Rate of the Public Log-Likelihood Ratio)
Let $f(n)$ be a function of $n$. We say that $f(n)$ characterizes the convergence rate of $l_{n}$ if, conditional on $\theta=+1$,
$\lim_{n\to\infty}\frac{l_{n}}{f(n)}=1\quad\text{with probability 1}.$
(2) Time to First Correct Action. This measure concerns how quickly the learning process produces a correct decision. The stopping time $\tau$ is defined as the index of the first agent whose action matches the true state:
Definition 4.4 (Time to First Correct Action)
The stopping time $\tau$ is given by
$\tau:=\inf\{n\geq 1:x_{n}=\theta\},$
where $x_{n}$ is the action of agent $n$. Learning is efficient under this metric if $\mathbb{E}[\tau]<\infty$.
(3) Total Number of Incorrect Actions. A stricter efficiency measure tracks the cumulative number of incorrect decisions across all agents:
Definition 4.5 (Total Number of Incorrect Actions)
Let $(x_{n})_{n\geq 1}$ be the sequence of actions and let $\theta$ be the true state. The total number of incorrect actions is defined as:
$W:=\sum_{n=1}^{\infty}\mathbf{1}\{x_{n}\neq\theta\}.$
Since $W\geq\tau-1$, this metric is at least as strict as the stopping time. We say that learning is efficient in this sense if $\mathbb{E}[W]<\infty$.
In the following sections, we analyze these three measures under privacy constraints and demonstrate that—perhaps counterintuitively—learning efficiency can be improved under privacy, compared to the non-private case. For example, Rosenberg and Vieille (2019) and Hann-Caruthers et al. (2018) show that in the Gaussian model without privacy protection, learning is asymptotically correct but highly inefficient. Specifically:
  • The convergence rate of the public belief is only $\Theta(\sqrt{\log n})$, much slower than the $\Theta(n)$ rate achievable under direct signal sharing Hann-Caruthers et al. (2018), Rosenberg and Vieille (2019);
  • The expected stopping time is infinite: $\mathbb{E}[\tau]=\infty$;
  • The expected number of incorrect actions is also infinite: $\mathbb{E}[W]=\infty$.
These results underscore the surprising inefficiency of the standard Gaussian sequential learning model. In contrast, we will show that when agents adopt certain privacy-preserving strategies, learning not only remains asymptotically correct but also becomes substantially more efficient under all three measures.
4.1 Main Results for Continuous Signals
For continuous signals, the sensitivity of an agent’s action to their private signal varies across the signal space. In particular, near the decision threshold, small changes in $s_{n}$ can flip the action from 1-1 to +1+1, whereas far from the threshold, even significant changes in $s_{n}$ may not alter the action at all. This heterogeneity in sensitivity highlights a key limitation of traditional differential privacy: under its uniform adjacency definition, it requires strategies to protect against worst-case signal perturbations, thereby injecting excessive noise even in regions where the action is insensitive. To address this, we adopt metric differential privacy, which allows the privacy guarantee to degrade gracefully with the distance between private signals. Under mDP, the strategy is required to produce similar outputs only for similar inputs, meaning that when two signals are far apart in value, the induced outputs (i.e., actions) can differ substantially without violating the privacy constraint. This localized notion of sensitivity enables us to reduce the noise injected in regions where the signal-to-action map is stable. Building on this observation, we propose the smooth randomized response strategy as the optimal strategy for agents under mDP. This strategy endogenously calibrates the amount of noise to the agent’s signal, injecting more noise near the decision threshold—where sensitivity is high—and less noise elsewhere. Inspired by the idea of smooth sensitivity (Nissim et al. 2007), this approach ensures that the strategy remains stable and metic differentially private even in settings with variable local sensitivity, while avoiding the inefficiencies of conventional randomized response. Crucially, the smooth randomized response strategy leverages the structure of mDP to strike a balance between privacy protection and information transmission. By tailoring the noise to the local geometry of the signal space, it allows agents to maintain stronger privacy guarantees with less informational distortion, thereby improving both privacy preservation and learning efficiency in sequential decision-making.
Definition 4.6 (Smooth Randomized Response)
A strategy $\mathcal{M}(s_{n};h_{n-1})$ is called a smooth randomized response strategy with flip probability $u_{n}$ if, given an initial decision $a_{n}$, the reported action $x_{n}$ is determined as follows:
$\mathbb{P}_{\mathcal{M}}(x_{n}=+1\mid a_{n}=-1)=\mathbb{P}_{\mathcal{M}}(x_{n}=-1\mid a_{n}=+1)=u_{n}(s_{n}),$
$\mathbb{P}_{\mathcal{M}}(x_{n}=+1\mid a_{n}=+1)=\mathbb{P}_{\mathcal{M}}(x_{n}=-1\mid a_{n}=-1)=1-u_{n}(s_{n}),$
with the flip probability $u_{n}$ defined as:
$u_{n}(s_{n})=\frac{1}{2}\,e^{-\varepsilon|s_{n}-t(l_{n})|}$
Here, $t(l_{n})=-\frac{\sigma^{2}l_{n}}{2}$ represents the threshold value for different actions as a function of the public belief $l_{n}$. If $s_{n}>t(l_{n})$, agents prefer to choose action +1+1 before flipping the action. Conversely, if $s_{n}<t(l_{n})$, agents prefer to choose action 1-1 before flipping.
Theorem 4.7 (Smooth Randomized Response Strategy)
In the sequential learning model with Gaussian signals, the optimal reporting strategy under a fixed privacy budget $\varepsilon$ is the smooth randomized response strategy. This strategy satisfies $(\varepsilon,d_{\mathbb{R}})$-metric differential privacy.
Unlike the randomized response, the smooth randomized response strategy adds noise based on the received signal, with the amount of noise decreasing as the distance from the threshold increases. For signals $s_{n}$ far from the threshold, agents add minimal noise, allowing them to preserve metric differential privacy without the need to add excessive noise, as required by the randomized response. This approach enables asymptotic learning to remain feasible even under privacy concerns because the reduction in noise allows agents to make more accurate decisions based on their signals.
Theorem 4.8 (Asymptotic Learning with Smooth Randomized Response)
For differentially private sequential learning with binary states and Gaussian signals, asymptotic learning will always occur under a smooth randomized response strategy with any privacy budget $\varepsilon\in(0,\infty)$.
Theorem 4.8 implies that, under the smooth randomized response strategy, agents will eventually take the correct action with a probability approaching one for any privacy budget $\varepsilon\in(0,\infty)$. In other words, despite privacy restrictions, the strategy allows agents to learn accurately over time, ensuring reliable decision making in the long run.
Having established the guarantee of asymptotic learning under the smooth randomized response, a natural follow-up question is whether the speed of asymptotic learning in the private regime can exceed the non-private baseline. Interestingly, our findings show this to be true. We measure the speed of learning by examining how public belief evolves as more agents act ($n\to\infty$). The convergence rate of the public log-likelihood ratio serves as an effective measure to quantify asymptotic learning, as it reflects the rate at which public beliefs converge to the correct state (Hann-Caruthers et al. 2018). Moreover, it forms the analytical basis for several other metrics, such as the expected time until the first correct action (Rosenberg and Vieille 2019) and the expected time until the correct cascade occurs (Hann-Caruthers et al. 2018).
Theorem 4.9 (Learning Rate under Smooth Randomized Response)
Consider a differential privacy budget $\varepsilon$ and a smooth randomized response strategy for sequential learning with Gaussian signals. Let $C(\varepsilon)=\frac{\varepsilon\sigma^{2}}{4}\left(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}\right)$. For $\varepsilon\in(0,\infty)$, the convergence rate of the public log-likelihood ratio under this strategy is $f(n)=\frac{2}{\varepsilon\sigma^{2}}\log(C(\varepsilon)n)\sim\frac{2}{\varepsilon\sigma^{2}}\log(n)$, where $f(n)\sim g(n)$ means $\lim_{n\to\infty}\frac{f(n)}{g(n)}=1$.
The $\Theta_{\varepsilon}(\log(n))$ asymptotic learning rate in Theorem 4.9 represents an improvement by an order of $\sqrt{\log(n)}$ over the $\Theta(\sqrt{\log(n)})$ asymptotic learning speed established by Hann-Caruthers et al. (2018) in the non-private regime. To better understand this improvement, assume that $\theta=+1$. For Gaussian signals, asymptotic learning is guaranteed, which implies that over time, more agents will take action +1+1 than 1-1. To protect their private signals, agents use the smooth randomized response to add noise to their actions. Importantly, this noise is asymmetric: In the true state, where signals are more aligned with the majority, agents require less distortion to maintain privacy; under the false state, where signals are more likely to contradict the majority, agents need to introduce more noise. This asymmetry arises because privacy protection requires obscuring how surprising a signal is — more surprising signals (those against the majority) require more distortion. As a result, actions under the correct state remain more stable, while those under the wrong state are noisier. This amplifies the difference between the two states in terms of observed behavior and makes actions more informative for later agents, thereby accelerating learning.
Figure 3: Smooth randomized response and the signal likelihoods. (a) shows the signal distribution under two different states ($\theta=-1$ in blue and $\theta=+1$ in orange), with the shaded area representing the probability of choosing $a_{n}=+1$ in each case. The vertical dotted line indicates the threshold location. (b) shows the smooth randomized response function as it varies with $s_{n}$ (black). (c) shows the probability of action changing from +1+1 to 1-1 under the influence of smooth randomized response, represented by the shaded area.
This intuition is further illustrated in Figure 3. Although the smooth randomized response reduces the probability of taking action +1+1 in both states, the reduction is more pronounced when $\theta=-1$, due to the shape of the signal distribution and its distance from the decision threshold. Specifically, the inferred flipping probability, i.e., the probability that a correct action is flipped to incorrect, as inferred by the next agent, is higher in the wrong state than in the true state. Because agent $n+1$ does not observe $s_{n}$, it must infer how agent $n$ might have acted on different possible signals using the smooth randomized response. This leads to the log-likelihood update:
$l_{n+1}=l_{n}+\log\frac{\mathbb{P}(x_{n}=+1\mid l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1\mid l_{n},\theta=-1)}.$ (6)
Both terms in the ratio decrease due to noise, but the denominator decreases more sharply. This leads to a growing log-likelihood ratio over time, making the signal in actions stronger and facilitating faster convergence.
In our framework, the proposed smooth randomized response strategy is in fact optimal in the sense that it achieves the largest asymmetry between the two states under $\varepsilon$-metric differential privacy constraints. Specifically, to guarantee $\varepsilon$-metic differential privacy, the ratio of flip probabilities for any pair of adjacent signals is bounded above by $e^{\varepsilon}$. The smooth randomized response strategy is constructed so that this ratio, which represents the rate at which the flipping probability decreases as the signal moves away from the decision threshold, exactly reaches the bound $e^{\varepsilon}$. This ensures that the strategy attains the maximal asymmetry in flipping behavior between the true and false states. Such asymmetry amplifies the difference in how actions are interpreted across states and allows actions to convey the strongest possible support for the true state under privacy constraints. As a result, the smooth randomized response not only satisfies differential privacy but also achieves the most efficient information aggregation enabled by this maximal asymmetry.
It is worth mentioning that the strategy we propose differs from existing studies. There are two distinct ways to accelerate learning speed. One approach is to modify the tail properties of the signal distribution. Rosenberg and Vieille (2019) and Hann-Caruthers et al. (2018) observe that Gaussian signals lead to slow learning speeds due to their thin tails, which reduce the probability of reversing an incorrect trend. This occurs because the likelihood of receiving a strong enough private signal to counteract an incorrect cascade is very low. In contrast, heavy-tailed distributions allow for rare but highly informative signals that can quickly correct mistaken beliefs. Another way to accelerate learning speed is to increase the weight of private signals. Arieli et al. (2025) show that overconfidence accelerates learning, as agents place greater reliance on their private signals rather than following public consensus. However, these approaches either depend on fixed signal properties or introduce systematic biases. In our paper, we leverage the asymmetry of the smooth randomized response, which selectively enhances the informativeness of correct actions while maintaining privacy constraints. Unlike a standard randomized response, which introduces noise indiscriminately, our strategy preserves useful information in decision making, leading to more efficient information aggregation and improved learning speed.
Having established that the convergence rate of public beliefs can be significantly improved under privacy-preserving strategy, we now turn our attention to the other two measures of learning efficiency: the expected time to the first correct action and the expected number of incorrect actions. While convergence rate captures long-run information aggregation, these two measures are particularly important for assessing how quickly correct behavior emerges and how many errors accumulate along the way. The following theorem shows that, under our proposed privacy-aware strategy, the expected stopping time for the first correct action may remain finite. This result confirms that learning under privacy constraints is not only asymptotically accurate but also timely in practice.
Theorem 4.10 (Learning Efficiency: Finite Expected Time to First Correct Action)
Consider a fixed privacy budget and suppose that agents follow a smooth randomized response strategy under Gaussian signals in a sequential learning setting. Then, the expected time $\mathbb{E}[\tau]$ until the first correct action satisfies
$\mathbb{E}[\tau]=C_{1}C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}},$ (7)
where $C_{1}$ is a positive constant that does not depend on $\varepsilon$. The series in Equation 7 converges if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$; and thus, the expected time to the first correct action is finite $(\mathbb{E}[\tau]<+\infty)$, if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$.
To provide insights into Theorem 4.10, we outline the proof idea here with detailed mathematical derivations provided in Section 13. We first define the public belief of agent $n$, denoted by $\pi_{n}$, as
$\pi_{n}=\mathbb{P}(\theta=+1\mid x_{1},x_{2},\dots,x_{n-1}),$
which evolves over time based on the sequence of previous decisions $x_{1},x_{2},\dots,x_{n-1}$.
We also define a special case $\pi_{n}^{*}$, representing the public belief of agent $n$ under full consensus, i.e., when all previous agents have reported +1+1:
$\pi_{n}^{*}=\mathbb{P}(\theta=+1\mid x_{1}=x_{2}=\dots=x_{n-1}=+1).$
In particular, the public belief in the next round, $\pi_{n+1}^{*}$, is derived from $\pi_{n}^{*}$ using the relationship given by Equation 6:
$\frac{\pi_{n+1}^{*}}{1-\pi_{n+1}^{*}}=\frac{\pi_{n}^{*}}{1-\pi_{n}^{*}}\times\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}.$ (8)
This equation describes the Bayesian update of the public belief as it transitions between rounds in the context of learning under privacy constraints. Assume $\theta=-1$. The probability $u_{n}:=P_{-}(\tau>n)$, representing the event that the first correct guess does not occur earlier than in round $n+1$, is given by:
$u_{n}=P_{-}(\tau>n)=P_{-}(x_{1}=\cdots=x_{n}=+1)=\prod_{k=1}^{n}\mathbb{P}(x_{k}=+1|l_{k},\theta=-1),$ (9)
where $\mathbb{P}(x_{k}=+1|l_{k},\theta=-1)$ represents the probability that the $k$-th guess is incorrect under the smooth randomized response strategy.
Since $\mathbb{E}_{-}(\tau)=\sum_{n=1}^{\infty}u_{n}$, the key question becomes whether the series $\sum_{n=1}^{\infty}u_{n}$ converges. This convergence guarantees that the expected stopping time $\mathbb{E}_{-}(\tau)$ is finite, ensuring efficient learning. Based on Equations 6, 8 and 9, we have
$u_{n}=\frac{1-\pi_{n+1}^{*}}{\pi_{n+1}^{*}}\prod_{k=1}^{n}\mathbb{P}(x_{k}=+1|l_{k},\theta=+1).$ (10)
A martingale argument shows that $\mathbb{P}(x_{k}=+1|l_{k},\theta=+1)$ is positive and $\pi_{n+1}^{*}\geq\frac{1}{2}$; therefore, we have $u_{n}\sim 1-\pi_{n+1}^{*}$. Since $l_{n}=\log\left(\frac{\pi_{n}^{*}}{1-\pi_{n}^{*}}\right)$, we have $1-\pi_{n}^{*}\sim e^{-l_{n}}$, where $l_{n}=\frac{2}{\varepsilon\sigma^{2}}\ln(C(\varepsilon)n)+c$ where $c$ is a constant. Thus, the expected time $\mathbb{E}[\tau]$ until the first correct action satisfies
$\mathbb{E}[\tau]=\sum_{n=1}^{\infty}u_{n}=C_{1}C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}},$
for some constant $C(\varepsilon)$ depending on $\varepsilon$. Since $\varepsilon<\frac{2}{\sigma^{2}}$, the series converges and thus the expected time until the first correct action is finite.
We show that under this measure of learning efficiency, the expected time for the first correct guess is finite when the privacy budget is limited. Interestingly, in the non-privacy setting, where the privacy budget is infinite, learning becomes inefficient, as the expected time for the first correct guess is infinity. This counterintuitive finding further supports our earlier conclusion that improving privacy protection can align with an improvement in learning speed. In fact, our results suggest that tighter privacy concerns do not necessarily lead to slower learning. The smooth randomized strategy internalizes the trade-off between revealing information and preserving privacy, leading to improved learning efficiency without violating privacy constraints. This insight challenges the traditional view of privacy and learning as conflicting objectives, and instead reveals how the two can be jointly optimized through endogenous adaptation.
In addition to the stopping time, we further show that the total number of incorrect actions may also have finite expectation under privacy concerns.
Theorem 4.11 (Learning Efficiency: Finite Expected Total Number of Incorrect Actions)
Consider a fixed privacy budget and suppose that agents follow a smooth randomized response strategy under Gaussian signals in a sequential learning setting. Then, the expected total number of incorrect actions $\mathbb{E}[W]$ satisfies
$\mathbb{E}[W]=C_{2}C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}},$ (11)
where $C_{2}$ is a positive constant that does not depend on $\varepsilon$. The series in Equation 11 converges if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$; and thus, the expected total number of incorrect actions is finite, $\mathbb{E}[W]<+\infty$, if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$.
To provide insights into Theorem 4.11, we outline the proof idea here, with all the details available in Section 14. First, we divide the action sequence $x_{n}$ into two parts: good runs and bad runs. A good run (respectively, a bad run) is a maximal sequence of consecutive actions that align (or fail to align) with the true state.
We then show that the duration of bad runs can be bounded by a geometric series, ensuring that the total sum remains finite. Specifically, we first prove that the conditional expected duration
$\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}|\mathcal{H}_{\tau_{k-1}}]\leq\frac{\pi_{\tau_{k-1}}}{1-\pi_{\tau_{k-1}}}\mathbb{E}[\tau]\leq\mathbb{E}[\tau]$
where $\tau_{k-1}$ is the public belief at the beginning of the (k1)(k-1)th good run, and $\mathcal{H}_{\tau_{k-1}}$ represents the corresponding $\sigma$-algebra. This follows from Theorem 4.10, which establishes that the expected time until the first correct action for the limited privacy budget is finite. Therefore, using this result, we further extend our proof to bound the unconditional expectation $\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}]$. We establish that this expectation follows an upper-bound recurrence relationship, where as $k$ increases, the expected duration of bad runs decreases. This occurs because the probability of a new bad run starting diminishes with increasing $k$, ensuring that the total number of incorrect actions remains finite.
Under a more stringent measure of learning efficiency, we find that learning in a privacy-preserving setting remains efficient as long as the privacy budget is limited. This result is particularly counterintuitive, as conventional wisdom suggests that the introduction of privacy-aware strategies inherently reduces the quality of information transmission. However, our findings reveal that learning can actually become more efficient due to improvements in the information aggregation process.
However, this efficiency gain is not without limitations. In particular, if the asymmetry becomes too extreme, such as when the privacy budget $\varepsilon$ becomes very small, the benefit may diminish. As $\varepsilon\to 0$, the scaling term $C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}$ diverges, potentially offsetting the gains in informativeness and leading to a deterioration in learning efficiency. Therefore, our theoretical guarantees assume that the ratio $\varepsilon$ is kept constant and bounded away from zero, ensuring that the overall effect of the randomized response remains beneficial.
When $\varepsilon<\frac{2}{\sigma^{2}}$, both the expected time to the first correct action and the expected total number of incorrect actions are finite. This raises a natural question: what is the optimal privacy budget $\varepsilon^{*}$ that minimizes these two expectations? Both quantities are determined by the expression $C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}}$, and our goal is to find the value of $\varepsilon^{*}$ that minimizes this term. However, due to the involvement of the Riemann zeta function in the infinite sum, obtaining a closed-form solution is infeasible. Therefore, we set $\sigma=\sqrt{2}$ so that $\tfrac{2}{\sigma^{2}}=1$, which normalizes the range of $\varepsilon$ to $(0,1)$. Accordingly, in Figure 4 we numerically obtain the optimal privacy budget $\varepsilon^{*}\approx 0.76$. While $\varepsilon$ is endogenously chosen by agents in equilibrium, this calculation serves as a normative benchmark for a planner or platform that can influence privacy choices through incentives or design. Such a benchmark helps quantify the potential welfare loss from purely self-interested privacy decisions and highlights the scope for policy interventions that align individual incentives with socially optimal information aggregation, echoing broader discussions on principled $\varepsilon$-selection in both practice (Dwork et al. 2019) and economics (Hsu et al. 2014).
Figure 4: The value of $C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}}$ as a function of the privacy budget $\varepsilon$ ($\sigma=\sqrt{2}$).
The Continuous Signals Model with Heterogeneous Privacy Budgets
In the homogeneous setting, we have already characterized the optimal privacy budget $\varepsilon^{*}$ for minimizing both the expected time to the first correct action and the expected total number of incorrect actions, assuming a fixed and positive privacy budget $\varepsilon$. However, the optimal convergence rate of learning in this setting remains unknown when $\varepsilon$ approaches zero. Interestingly, as $\varepsilon$ approaches zero, the speed of learning can increase. To investigate this phenomenon, we consider a heterogeneous setting in which each agent’s privacy budget $\varepsilon_{n}$ is drawn from a distribution that assigns positive probability mass to values of $\varepsilon$ close to zero. This captures agents with varying privacy concerns. We assume $\varepsilon_{n}\sim U[0,1]$ for simplicity, noting that the results extend to $\varepsilon_{n}\sim U[0,a]$ for any $a>0$ after appropriate rescaling. Additional simulation results for different choices of privacy budget intervals and their corresponding learning performance are provided in Figure 5. We show that the convergence rate of the public log-likelihood ratio can increase to $\Theta(\sqrt{n})$ due to the presence of agents with strong privacy preferences. Before presenting the formal results, we first establish that agents with $\varepsilon_{n}\sim U[0,1]$ still achieve asymptotic learning, as shown in Theorem 4.12.
Theorem 4.12 (Asymptotic Learning under Heterogeneous Privacy Budgets)
Consider a sequential learning model with binary states, Gaussian private signals, and agent-specific privacy budgets $\varepsilon_{n}$ drawn independently from a Uniform distribution $\mathrm{U}[0,1]$. Then, under a smooth randomized response strategy, asymptotic learning occurs almost surely.
Theorem 4.12 shows that asymptotic learning still occurs almost surely even when agents have heterogeneous privacy preferences, modeled by privacy budgets $\varepsilon_{n}$ drawn from a continuous uniform distribution $\mathrm{U}[0,1]$. This result confirms that the smooth randomized response strategy preserves the long-run accuracy of social learning despite the presence of highly privacy-conscious agents. In the next theorem, we go beyond asymptotic learning and characterize the speed at which public belief converges. Specifically, we quantify how the distribution of privacy budgets affects the convergence rate of the public log-likelihood ratio in this heterogeneous setting.
Theorem 4.13 (Learning Rate under Heterogeneous Privacy Budgets)
Consider differential privacy budget $\varepsilon_{n}$ follows $U[0,1]$ and a smooth randomized response strategy for sequential learning with Gaussian signals. Then, the convergence rate of the public log-likelihood ratio under this strategy is given by $f(n)=\Theta(\sqrt{n})$.
Due to the presence of highly privacy-conscious agents, the convergence rate of the public belief increases to $\Theta(\sqrt{n})$, which is significantly faster than the $\Theta(\log n)$ rate observed in the homogeneous setting. This naturally raises the question: is it possible to achieve an even faster convergence rate under some alternative distribution of privacy budgets? The following theorem answers this question by establishing a fundamental limit: regardless of the distribution from which agents’ privacy budgets are drawn, the convergence rate of the public log-likelihood ratio cannot exceed $\Theta(\sqrt{n})$.
Theorem 4.14 (Learning Rate Bound under Heterogeneous Privacy Budgets)
Consider a sequential learning model with Gaussian private signals and a smooth randomized response strategy, where each agent’s privacy budget $\varepsilon_{n}$ is drawn independently from a distribution $G$ supported on $[0,a]$. Suppose that $G$ assigns zero measure to the singleton $\{0\}$, i.e., $G(\{0\})=0$. Then, the convergence rate of the public log-likelihood ratio under this strategy is upper bounded by $f(n)=\Theta(\sqrt{n})$ and lower bounded by $f(n)=\Theta(\log n)$.
Theorem 4.14 highlights a fundamental limit in privacy-aware sequential learning: the presence of agents with very small privacy budgets effectively determines the maximal rate at which public belief can converge. To place this result in context, it is useful to compare it with the following two benchmarks. In the ideal case where an agent directly observe $n$ i.i.d. Gaussian private signals, the convergence rate of the aggregated belief is $\Theta(n)$, corresponding to the classical law of large numbers. In the standard non-private sequential learning setting (i.e., without any privacy-preserving behavior), the convergence rate is reduced to $\Theta(\sqrt{\log(n)})$ due to the information loss from agents observing only the actions of their predecessors (Hann-Caruthers et al. 2018, Rosenberg and Vieille 2019).
Our result shows that, when smooth randomized response is employed under heterogeneous privacy budgets, the convergence rate can improve to $\Theta(\sqrt{n})$, which is strictly faster than that in the non-private sequential setting. This somewhat surprising outcome arises from two key factors: the presence of highly privacy-conscious agents (with small $\varepsilon_{n}$) and the inherent asymmetry of the smooth randomized response strategy under the two possible states. Specifically, under the true state, the average probability of flipping the action is lower than under the false state, which makes the observed actions more informative in expectation. Nonetheless, Theorem 4.14 confirms that $\Theta(\sqrt{n})$ is the fastest achievable convergence rate under any distribution of privacy budgets, and thus represents the optimal learning rate under privacy constraints in sequential settings.
To further support our results in the heterogeneous setting, we also analyze two important measures of learning efficiency: the expected time until the first correct action, and the expected total number of incorrect actions. We find that both quantities remain finite under the smooth randomized response strategy with heterogeneous privacy budgets, which highlights that privacy-aware learning can still be highly efficient even when agents have varying degrees of privacy concerns.
Theorem 4.15 (Time to First Correct Action under Heterogeneous Privacy Budgets)
Consider a privacy budget $\varepsilon_{n}$ follows follows $U[0,1]$ and suppose that agents follow a smooth randomized response strategy under Gaussian signals in a sequential learning setting. Then, the expected time $\mathbb{E}[\tau]$ until the first correct action satisfies
$\mathbb{E}[\tau]=C_{1}\sum_{n=1}^{\infty}e^{-\tilde{C}n^{\frac{1}{2}}},$ (12)
where $C_{1}$ and $\tilde{C}$ are positive constants . The series in Equation 12 is finite.
Theorem 4.16 (Total Number of Incorrect Actions under Heterogeneous Privacy Budgets)
Consider a privacy budget $\varepsilon_{n}$ follows follows $U[0,1]$ and suppose that agents follow a smooth randomized response strategy under Gaussian signals in a sequential learning setting. Then, the expected total number of incorrect actions $\mathbb{E}[W]$ satisfies
$\mathbb{E}[W]=C_{2}\sum_{n=1}^{\infty}e^{-\tilde{C}n^{\frac{1}{2}}},$ (13)
where $C_{2}$ and $\tilde{C}$ are positive constants. The series in Equation 13 is finite.
Theorems 4.15 and 4.16 together demonstrate that, even in the presence of heterogeneous privacy budgets, the learning process remains efficient. Specifically, agents are expected to take a correct action in finite time, and the overall number of incorrect actions remains bounded. These results further confirm that smooth randomized response enables timely and reliable decision-making under privacy constraints.
Simulation under Heterogeneous Privacy Budgets.
While our theory shows that stronger privacy concerns yield faster asymptotic learning (as $n\to\infty$), we ask how learning behaves in finite populations. Following Alaggan et al. (2015), Acharya et al. (2025), we classify agents by privacy: conservatives ($\varepsilon=0.1$), pragmatists ($\varepsilon=0.5$), and liberals ($\varepsilon=1$). We study five regimes: three homogeneous settings with fixed $\varepsilon\in{0.1,0.5,1}$, a heterogeneous setting with $\varepsilon\sim U(0,1)$, and a non-private baseline. For each regime we run five independent simulations; signals are normal with standard deviation $\sigma=1$. This design ties behavioral types to budgets and tests whether asymptotic predictions already manifest in finite samples.
Figure 5: Log-likelihood ratio dynamics under five privacy regimes. Homogeneous settings: conservatives $(\varepsilon=0.1)$, pragmatists $(\varepsilon=0.5)$, and liberals $(\varepsilon=1)$; heterogeneous setting: $\varepsilon\sim U(0,1)$; and a non-private baseline. Signals are normally distributed with $\sigma=1$.
Figure 5 illustrates the trajectory of the log-likelihood ratio $l_{n}$ as the number of agents increases. We observe that, consistent with the theoretical result, the log-likelihood ratio grows more rapidly under stronger privacy concerns (lower budgets). Intuitively, when privacy concern is stronger, agents introduce more randomization in their actions. This randomization increases the gap between action distributions under the two states: under the true state, agents conform to the majority action more reliably, whereas under the false state, their actions are noisier and deviate more often. This asymmetry amplifies the log-likelihood ratio between states, thereby accelerating information aggregation.
5 Conclusions
We investigated the effects of privacy constraints on sequential learning, examining binary and continuous signals under a metric differential privacy framework. Our results reveal that in the binary signal case, privacy-preserving noise impairs learning performance by reducing the informativeness of individual actions. Interestingly, the probability of achieving the correct cascade in the binary case is not monotonic with respect to the privacy budget $\varepsilon$. Specifically, there are breakpoints at the endpoints of intervals where increasing the privacy budget can paradoxically decrease the probability of the correct cascade, driven by changes in the information cascade threshold. These findings highlight the challenges posed by binary signals in achieving efficient learning under privacy constraints.
For continuous signals, privacy-aware agents adopt the smooth randomized response strategy, which adjusts noise based on the signal’s distance from the decision threshold. This ensures metric differential privacy while reducing unnecessary noise for signals and amplifying informativeness differences between actions. The strategy introduces asymmetry that makes correct actions more informative, enhancing belief updates and accelerating learning. As a result, asymptotic learning is not only preserved under privacy but also occurs at an improved rate of $\Theta_{\varepsilon}(\log n)$, compared to $\Theta(\sqrt{\log n})$ without privacy. These findings highlight how adaptive noise strategies can improve learning efficiency even under stringent privacy constraints.
We show that under the smooth randomized response, the expected time to the first correct action, $\mathbb{E}[\tau]$, is finite when $\varepsilon\in(0,\frac{2}{\sigma^{2}})$, ensuring efficient learning. We also introduce a stronger metric, the expected number of incorrect actions, $\mathbb{E}[W]$, and prove it remains finite under privacy. In contrast, both quantities diverge in the non-private regime. These results challenge the traditional trade-off between privacy and learning, showing that strong privacy does not necessarily slow down learning. Instead, adaptive noise strategies like the smooth randomized response can simultaneously preserve privacy and enhance learning efficiency in sequential settings with continuous signals.
We further consider a heterogeneous setting where agents draw their privacy budgets from a uniform distribution. Interestingly, the presence of agents with strong privacy preferences—those with smaller budgets—leads to an acceleration in the aggregation of public belief, achieving a convergence rate of order $\Theta(\sqrt{n})$. This rate is shown to be the fastest possible under any distribution of privacy budgets. These findings highlight two key points: first, carefully structured privacy constraints can paradoxically enhance learning speed; and second, despite this improvement, sequential decision-making under privacy remains fundamentally limited in its ability to match the efficiency of fully transparent environments, where direct access to signals enables linear convergence.
Our findings offer practical insights for organizations balancing user privacy with accurate decision-making. Contrary to the belief that privacy hinders learning, we show that adaptive noise strategies—like smooth randomized response—can both protect privacy and accelerate learning. This is valuable for platforms aggregating consumer or health data, where timely and accurate insights matter. Moreover, the advantages of heterogeneous privacy preferences suggest that encouraging voluntary, varied participation may outperform rigid disclosure policies. Overall, our results provide guidance for designing privacy-aware systems that uphold ethical standards while supporting organizational goals in learning and adaptation.
[…]
We thank Juba Ziani, Krishna Dasaratha, David Hirshleifer and Jie Gao for their valuable feedback and discussions.
Code Availability
All R scripts used to generate the figures and reproduce the simulation results in this paper are openly available on our GitHub repository: https://github.com/YuxinLiu1997/Privacy-Preserving-Sequential-Learning.
Appendix
6 Additional Related Work
Social learning under model misspecification.
Our research relates to the literature on social learning under model misspecification, which typically emphasizes the adverse consequences of imperfect models on learning outcomes. For example, Frick et al. (2023) show that even minor misspecifications can substantially hinder belief convergence, while Le et al. (2017) demonstrate that exogenous observation errors increase the likelihood of incorrect cascades as noise grows. Similarly, Bohren (2016) highlight how mistaken assumptions about correlation structures among others’ actions can generate persistent inefficiencies. These studies collectively suggest that misspecification erodes the reliability of information aggregation in sequential environments. In contrast, we show that when privacy concerns endogenously induce agents to randomize their actions, the resulting noise in the continuous-signal setting can paradoxically enhance asymptotic learning by amplifying informative action and accelerating convergence to the truth.
Privacy concern and information disclosure.
A central theme in the privacy literature concerns how individuals make disclosure decisions when facing privacy constraints. Adjerid et al. (2019) demonstrate that choice architecture and framing effects can lead to cascaded patterns of privacy disclosure, suggesting that even subtle interventions systematically shift disclosure behaviors. Ke and Sudhir (2023) study how regulatory regimes such as GDPR reshape individual incentives and the structure of personal data markets, while Kim et al. (2025) highlight the autonomy paradox that arises when people consent to behavior-based pricing, balancing perceived control with potential welfare losses. Similarly, Casadesus-Masanell and Hervas-Drane (2015) and Montes et al. (2019) analyze how the value of personal information and endogenous privacy preferences shape disclosure incentives in market settings. These studies establish that privacy fundamentally alters disclosure behavior, but they largely focus on individual decision-making or static market equilibria. In contrast, our work emphasizes the collective dimension: when agents in sequential learning environments strategically withhold or randomize disclosures under privacy concerns, the dynamics of inference and belief updating themselves are transformed. By analyzing correlated disclosure across agents, we show how privacy generates systematic effects on the speed and efficiency of information aggregation beyond the scope of individual-level disclosure choices.
Privacy protection and utility tradeoffs.
Our work also relates to the broader literature on the trade-off between privacy guarantees and utility, particularly in the context of relaxed definitions of differential privacy. Standard DP, while widely adopted, often suffers from limitations in settings with continuous or unbounded domains, as it treats all neighboring datasets as equally distinct and thus requires injecting excessive noise, leading to severe utility loss (Dwork and Roth 2014, Bun and Steinke 2016). To address these issues, several relaxed notions have been proposed. Metric DP (mDP) generalizes DP by replacing binary adjacency with distance-based similarity, allowing noise magnitude to scale with input proximity and enabling finer privacy–utility tradeoffs in applications such as geo-location, speech, and image data (Andrés et al. 2013, Han et al. 2020, Fan 2019). Pufferfish privacy provides another flexible framework that explicitly specifies which secrets should be indistinguishable, thereby avoiding unnecessary noise (Kifer and Machanavajjhala 2014); variants such as Blowfish further demonstrate improved utility for tasks like histograms and clustering (He et al. 2014, Song et al. 2017). Building on this line of work, we apply the metric differential privacy framework to study how privacy constraints affect sequential learning with continuous signals, and show how it achieves a balance between utility and privacy guarantees.
7 Proof of Theorem 3.4: Randomized Response Strategy for Binary Model
Before the information cascade, agents take actions entirely based on their private signals. Thus, the probability expressions become:
[…]
Now, using Equation 2, we can rewrite:
$\frac{\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})}{\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n})}$ $=\frac{\mathbb{P}(x_{n}\mid s_{n},x_{1},x_{2},\dots,x_{n-1})}{\mathbb{P}(x_{n}\mid s_{n}^{\prime},x_{1},x_{2},\dots,x_{n-1})}$
$\leq\max_{\begin{subarray}{c}x_{n}\in\{-1,+1\}\\ s_{n},\tilde{s}_{n}\in\{-1,+1\}\end{subarray}}\frac{\mathbb{P}(x_{n}\mid s_{n},x_{1},x_{2},\dots,x_{n-1})}{\mathbb{P}(x_{n}\mid\tilde{s}_{n},x_{1},x_{2},\dots,x_{n-1})}$
$=\max_{\begin{subarray}{c}x_{n}\in\{-1,+1\}\\ s_{n},\tilde{s}_{n}\in\{-1,+1\}\end{subarray}}\frac{\mathbb{P}_{\mathcal{M}_{n}}(x_{n}\mid a_{n}=+1)\mathbb{I}\{s_{n}=+1\}+\mathbb{P}_{\mathcal{M}_{n}}(x_{n}\mid a_{n}=-1)\mathbb{I}\{s_{n}=-1\}}{\mathbb{P}_{\mathcal{M}_{n}}(x_{n}\mid a_{n}=+1)\mathbb{I}\{\tilde{s}_{n}=+1\}+\mathbb{P}_{\mathcal{M}_{n}}(x_{n}\mid a_{n}=-1)\mathbb{I}\{\tilde{s}_{n}=-1\}}$
$=\frac{1-u_{n}}{u_{n}}=\exp(\varepsilon),$
After the information cascade, agents base their decisions solely on public history rather than their private signals. As a result, randomized response mechanisms for protecting private signals become unnecessary because private signals are no longer used in decision making. Finally, given the binary action space, the randomized response mechanism introduces the minimal perturbation required to satisfy the privacy constraint. As such, it is an optimal strategy among all mechanisms that preserve privacy.
8 Proof of Theorem 3.6: Probability of the Correct Cascade
Building on our observations in Section 7, we analyze the threshold of the information cascade. Here, we define $k$ as the difference between the occurrences of action +1+1 and action 1-1 in the observation history up to the agent $n$. With this in mind, we establish the following lemma.
Lemma 8.1
If the information cascade starts, the threshold $k$ is given by:
$k=\lfloor\log_{\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}}\frac{1-p}{p}\rfloor+1.$ (14)
Proof.
First, since the prior of the state is symmetric, the log-likelihood ratio of the public belief for agent $n$ based on the reports from agent $1$ to agent n1n-1 is defined in Equation 5:
$l_{n}=\log\frac{\mathbb{P}(\theta=+1|x_{1},\ldots,x_{n-1})}{\mathbb{P}(\theta=-1|x_{1},\ldots,x_{n-1})}=\log\frac{\mathbb{P}(x_{1},\ldots,x_{n-1}|\theta=+1)}{\mathbb{P}(x_{1},\ldots,x_{n-1}|\theta=-1)}.$
We begin the analysis by scrutinizing the action of agent 3. Suppose that before agent 3 takes action, the information cascade has not yet begun, and we examine the log-likelihood ratio of public belief for agent 3. Specifically, if $x_{1}=x_{2}=+1$, then,
$l_{3}=\log\frac{\mathbb{P}(x_{1}=+1|\theta=+1)}{\mathbb{P}(x_{1}=+1|\theta=-1)}\frac{\mathbb{P}(x_{2}=+1|\theta=+1,x_{1}=+1)}{\mathbb{P}(x_{2}=+1|\theta=-1,x_{1}=+1)}=\log\left[\frac{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}\right]^{2}.$
If $x_{1}=x_{2}=-1$, then
$l_{3}=\log\frac{\mathbb{P}(x_{1}=-1|\theta=+1)}{\mathbb{P}(x_{1}=-1|\theta=-1)}\frac{\mathbb{P}(x_{2}=-1|\theta=+1,x_{1}=-1)}{\mathbb{P}(x_{2}=-1|\theta=+1,x_{1}=-1)}=\log\left[\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}\right]^{2}.$
If $x_{1}=+1,x_{2}=-1$ or $x_{1}=-1,x_{2}=+1$, then $l_{3}=1$. For agent 3, in making decisions, it needs to compare the public log-likelihood ratio with the private log-likelihood ratio $L_{3}$.
$L_{3}=\log{\frac{\mathbb{P}(\theta=+1|s_{3})}{\mathbb{P}(\theta=-1|s_{3})}}=\log{\frac{\mathbb{P}(s_{3}|\theta=+1)}{\mathbb{P}(s_{3}|\theta=-1)}}.$
If $s_{3}=+1$, then $L_{3}=\log\frac{p}{1-p}$; otherwise, $L_{3}=\log\frac{1-p}{p}$. In equilibrium, agent 3 selects $a_{3}=+1$ if and only if $l_{3}+L_{3}>0$. If an information cascade occurs, agent 3 will choose action 1-1 after observing $x_{1}=x_{2}=-1$ even if $s_{3}=+1$, or agent 3 will choose action +1 after observing $x_{1}=x_{2}=+1$ even if $s_{3}=-1$. This implies the following expression.
$\log\left[\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}\right]^{2}+\log\frac{p}{1-p}<0,\log\left[\frac{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}\right]^{2}+\log\frac{1-p}{p}>0.$
For agent $n$, by induction we find that if the information cascade occurs, then
[…]
In total, if $k$ is the threshold for the start of the information cascade, then $k=\lfloor\log_{\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}}\frac{1-p}{p}\rfloor+1$.
Lemma 8.2
When $\tilde{u}(\varepsilon)\neq\frac{1}{2}$, the probability of a correct cascade is given by:
$\mathbb{P}(\text{correct cascade})=\frac{\rho(\varepsilon)^{k}-1}{\rho(\varepsilon)^{2k}-1}.$
where $\rho(\varepsilon)=\frac{1-\tilde{u}(\varepsilon)}{\tilde{u}(\varepsilon)}$ and $\tilde{u}(\varepsilon)=u(\varepsilon)(1-p)+p(1-u(\varepsilon))$. When $\tilde{u}(\varepsilon)=\frac{1}{2}$, the information cascade does not occur.
Proof.
For the case $\tilde{u}(\varepsilon)=\frac{1}{2}$, it is trivial. However, for $\tilde{u}(\varepsilon)\neq\frac{1}{2}$, we can reframe the problem as a Markov chain featuring two absorbing states: $k$ and k-k. The transition probabilities are defined as follows: $\mathbb{P}(z_{n}=m+1|z_{n-1}=m)=\tilde{u}(\varepsilon)$ and $\mathbb{P}(z_{n}=m-1|z_{n-1}=m)=1-\tilde{u}(\varepsilon)$, where $z_{n-1}$ is the difference between the occurrences of action +1+1 and action 1-1 in the observation history up to agent $n$. This formulation aligns with the classical gambler’s ruin problem.
Given an initial capital of $k$ defined in Equation 14, our objective is to calculate the probabilities associated with losing all capital $k$ or achieving a goal of $2k$. Without loss of generality, we assume $\theta=+1$, then the probability of achieving a goal of $2k$ is the same as the probability of the correct cascade. Let,
$S_{n}=X_{1}+\cdots+X_{n},\quad S_{0}=0.$
The event
$M_{k,n}=\left[k+S_{n}=2k\right]\cap\bigcap_{m=1}^{n-1}\left[0<k+S_{m}<2k\right],$
represents earning $2k$ at time $n$. If $s_{2k}(k)$ denotes the probability of achieving this goal ultimately, then
$s_{2k}(k)=P\left(\bigcup_{n=1}^{\infty}M_{k,n}\right)=\sum_{n=1}^{\infty}P\left(M_{k,n}\right).$
By convention, given the sample space $\Omega$ and the empty set $\emptyset$, we set $M_{k,0}=\emptyset$, $M_{2k,0}=\Omega$, and $M_{2k,n}=\emptyset$. Then we get $s_{2k}(0)=0$ and $s_{2k}(2k)=1$. Based on the problem setting, we could derive the following difference equation,
$s_{2k}(k)=\tilde{u}(\varepsilon)s_{2k}(k+1)+(1-\tilde{u}(\varepsilon))s_{2k}(k-1).$
By solving the above difference equation with side conditions $s_{2k}(0)=0$ and $s_{2k}(2k)=1$. We get the probability that the gambler can successfully achieve his goal before ruin as follows:
$s_{2k}(k)=\frac{\rho(\varepsilon)^{k}-1}{\rho(\varepsilon)^{2k}-1}.$
where $\rho(\varepsilon)=\frac{1-\tilde{u}(\varepsilon)}{\tilde{u}(\varepsilon)}$ and $\tilde{u}(\varepsilon)=u(\varepsilon)(1-p)+p(1-u(\varepsilon))$. ∎
Based on the above lemmas, we now present the formal proof of Theorem 4.11. First, we derive the interval $I_{k}$ in which the threshold is $k$. By Lemma 8.1, after the simple calculation, for any $k\geq 2$
$v_{k}=\frac{1-\alpha^{\frac{k-2}{k-1}}}{1-\alpha^{\frac{k-2}{k-1}}+\alpha^{\frac{-1}{k-1}}-\alpha},$
where $\alpha=(1-p)/p$. Thus, $I_{k}=(\varepsilon_{k+1},\varepsilon_{k}]$, and $\varepsilon_{k}=\log\frac{1-v_{k}}{v_{k}}$. By lemma 8.2 take the first derivative of $s_{2k}(k)$ with respect to $\rho(\varepsilon)$ for $k\geq 2$ and $\rho(\varepsilon)\in(0,1)$:
$\frac{ds_{2k}(k)}{d\rho}=\frac{3k\rho(\varepsilon)^{3k-2}-k\rho(\varepsilon)^{k-1}-2k\rho(\varepsilon)^{2k-1}}{(\rho(\varepsilon)^{2k}-1)^{2}}<0.$
Next, we take the first derivative of $\rho(\varepsilon)$ and $u(\varepsilon)$ with respect to $u$ and $\varepsilon$, respectively. Then $\frac{d\rho(\varepsilon)}{du(\varepsilon)}>0$ and $\frac{du(\varepsilon)}{d\varepsilon}<0$. Consequently. $\frac{ds_{2k}(k)}{d\varepsilon}>0$, With an increase in the differential privacy budget, the probability of a correct cascade also increases.
9 Failure of Asymptotic Learning for $\varepsilon$-Differential Privacy
In this appendix we show that under an $\varepsilon$-differential privacy constraint, agents adopt a randomized response strategy with a fixed flip probability, and consequently asymptotic learning fails to occur in the sequential learning model with Gaussian signals.
Theorem 9.1 (Randomized Response Strategy for Continuous Model)
In the sequential learning model with Gaussian signals under an $\varepsilon$-differential privacy constraint, agents adopt a randomized response strategy with a fixed flip probability
$u_{n}=u(\varepsilon)=\frac{1}{1+e^{\varepsilon}},$
independent of the agent index $n$.
Proof.
Proof of Proposition 9.1 we can derive the decision for agent $n$ before adding any noise. A simple calculation shows that agent $n$ takes action +1+1 iff $l_{n}+L_{n}>0$ which implies
$\log{\frac{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}|\theta=+1)}{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}|\theta=-1)}}+\log{\frac{\mathbb{P}(\theta=+1|s_{n})}{\mathbb{P}(\theta=-1|s_{n})}}>0.$
For Gaussian distribution
$L_{n}=\log{\frac{\mathbb{P}(\theta=+1|s_{n})}{\mathbb{P}(\theta=-1|s_{n})}}=\log\frac{f(s_{n}|\theta=+1)}{f(s_{n}|\theta=-1)}=\log{\frac{e^{-(s_{n}-1)^{2}/2\sigma^{2}}}{e^{-(s_{n}+1)^{2}/2\sigma^{2}}}}=\frac{2s_{n}}{\sigma^{2}}.$
That implies $L_{n}$ also follows a normal distribution. Let $G_{+}$ and $G_{-}$ denote the cumulative distribution functions of a Gaussian distribution with mean $1$ and variance $4$, respectively. Then, when $s_{n}>\frac{-l_{n}}{2}$, agent $n$ chooses action $a_{n}=+1$; otherwise, $a_{n}=-1$. This implies that $\mathbb{P}(a_{n}|s_{n},x_{1},x_{2},\ldots,x_{n-1})$ is either one or zero. Based on this observation and Equation 4, we could derive the relationship between the differential privacy budge $\varepsilon$ and the probability of randomized response $u_{n}$
$\hskip 12.0pt\frac{\mathbb{P}(\mathcal{M}(s_{n})=x_{n};h_{n-1})}{\mathbb{P}(\mathcal{M}(s_{n}^{\prime})=x_{n};h_{n-1})}$
$=\frac{\mathbb{P}(x_{n}|s_{n},x_{1},x_{2},\ldots,x_{n-1})}{\mathbb{P}(x_{n}|s_{n}^{\prime},x_{1},x_{2},\ldots,x_{n-1})}$
[…]
$=\frac{1-u_{n}}{u_{n}}=\exp(\varepsilon),$
where $\tilde{a_{n}}=-a_{n}$, then $u_{n}=\frac{1}{1+e^{\varepsilon}}$. ∎
Because each agent uses the same randomized strategy, even when their private signal provides strong evidence about the true state, the agent must still add noise with the same probability. This leads to frequent incorrect actions, even in highly informative scenarios, and gives rise to the following impossibility result:
Theorem 9.2 (Failure of Asymptotic Learning)
For $\varepsilon$-differentially private sequential learning with binary states and Gaussian signals, asymptotic learning does not occur under the randomized response strategy.
Proof.
Proof of Proposition 9.2 Without loss of generality, we assume $\theta=-1$, then the likelihood ratio of the public belief
$\pi_{n}:=\frac{\mathbb{P}(\theta=+1|x_{1},x_{2},\ldots,x_{n-1})}{\mathbb{P}(\theta=-1|x_{1},x_{2},\ldots,x_{n-1})},$
is a martingale conditional on $\theta=-1$. To see this,
$\pi_{n+1}$ $=\frac{\mathbb{P}(\theta=+1|x_{1},x_{2},\ldots,x_{n-1})}{\mathbb{P}(\theta=-1|x_{1},x_{2},\ldots,x_{n-1})}$
$=\frac{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}|\theta=+1)}{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}\theta=-1)}$
$=\pi_{n}\frac{\mathbb{P}(x_{n}|\pi_{n},\theta=+1)}{\mathbb{P}(x_{n}|\pi_{n},\theta=-1)}.$
Therefore, we have
$E(\pi_{n+1}|\pi_{n},\theta=-1)=\pi_{n}\sum_{x_{n}\in\{0,1\}}\frac{\mathbb{P}(x_{n}|\pi_{n},\theta=+1)}{\mathbb{P}(x_{n}|\pi_{n},\theta=-1)}\mathbb{P}(x_{n}|\pi_{n},\theta=-1)=\pi_{n}.$
Due to $\pi_{n}$ being a non-negative martingale, by martingale convergence theorem, with probability one, the limit $\pi_{\infty}=\lim_{n\to\infty}\pi_{n}$ exists and is finite, thus the limit $l_{\infty}=\lim_{n\to\infty}l_{n}$ also exists and is finite. Combined with the dynamics of the public log-likelihood ratio shown in Equation 6, $l_{\infty}$ has to satisfy $(1-u)G_{+}(-l_{n})+u(1-G_{+}(-l_{n}))=(1-u)G_{-}(-l_{n})+u(1-G_{-}(-l_{n}))$, then $\pi_{\infty}\in\{0,\infty\}$. In addition, $\pi_{\infty}<\infty$ implies $\pi_{n}=0$, $l_{\infty}=$-$\infty$.
$\lim_{n\to\infty}\mathbb{P}_{\mathcal{M}_{n}}(x_{n}=\theta)$
$=(1-u(\varepsilon))G_{-}(-l_{\infty})+u(\varepsilon)(1-G_{-}(-l_{\infty}))$
$=1-u(\varepsilon).$
In all, because agents always randomize actions for the purpose of protecting privacy with constant probability, asymptotic learning will not occur. ∎
10 Proof of Theorem 4.7: Smooth Randomized Response Strategy
To obtain an upper bound for the smooth randomized response strategy, we analyze the evolution of the public log-likelihood ratio. Without privacy concerns, agent $n$ chooses $a_{n}=+1$ if and only if
$l_{n}+L_{n}>0.$
Thus, there exists a threshold $t(l_{n})$ such that the agent chooses $a_{n}=+1$ if $s_{n}>t(l_{n})$, and $a_{n}=-1$ otherwise.
Suppose an agent receives a signal $s_{n}>t(l_{n})$, and another signal $s_{n}^{\prime}>t(l_{n})$ is also above the threshold. Then, by the smooth randomized response mechanism, we have
$\frac{\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})}{\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n})}\leq e^{\varepsilon d_{\mathbb{R}}(s_{n},s_{n}^{\prime})}.$
Now consider the case where $s_{n}>t(l_{n})$ but $s_{n}^{\prime}<t(l_{n})$. Then,
$\frac{\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})}{\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n})}=\max_{\begin{subarray}{c}x_{n}\in\{-1,+1\}\\ s_{n}>t(l_{n}),\;s_{n}^{\prime}<t(l_{n})\end{subarray}}\frac{\mathbb{P}(x_{n}\mid s_{n},h_{n-1})}{\mathbb{P}(x_{n}\mid s_{n}^{\prime},h_{n-1})}.$
If $x_{n}=+1$, then
$\frac{\mathbb{P}(x_{n}=+1\mid s_{n},h_{n-1})}{\mathbb{P}(x_{n}=+1\mid s_{n}^{\prime},h_{n-1})}$ $=\frac{1-\frac{1}{2}e^{-\varepsilon(s_{n}-t(l_{n}))}}{\frac{1}{2}e^{\varepsilon(s_{n}^{\prime}-t(l_{n}))}}$
$=e^{\varepsilon(s_{n}-s_{n}^{\prime})}\left(2e^{\varepsilon t(l_{n})-\varepsilon s_{n}}-e^{2\varepsilon t(l_{n})-2\varepsilon s_{n}}\right)$
$\leq e^{\varepsilon(s_{n}-s_{n}^{\prime})}.$
If $x_{n}=-1$, then
$\frac{\mathbb{P}(x_{n}=-1\mid s_{n},h_{n-1})}{\mathbb{P}(x_{n}=-1\mid s_{n}^{\prime},h_{n-1})}$ $=\frac{\frac{1}{2}e^{-\varepsilon(s_{n}^{\prime}-t(l_{n}))}}{1-\frac{1}{2}e^{\varepsilon(s_{n}-t(l_{n}))}}$
$=e^{\varepsilon(s_{n}-s_{n}^{\prime})}\cdot\frac{e^{-2\varepsilon s_{n}+\varepsilon t(l_{n})+\varepsilon s_{n}^{\prime}}}{2-e^{\varepsilon(s_{n}^{\prime}+t(l_{n}))}}$
$\leq e^{\varepsilon(s_{n}-s_{n}^{\prime})}.$
Now suppose $s_{n}<t(l_{n})$ and $s_{n}^{\prime}<t(l_{n})$. Then, it is straightforward that
$\frac{\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})}{\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n})}\leq e^{\varepsilon d_{\mathbb{R}}(s_{n},s_{n}^{\prime})}.$
If $x_{n}=+1$, we have
$\frac{\mathbb{P}(x_{n}=+1\mid s_{n},h_{n-1})}{\mathbb{P}(x_{n}=+1\mid s_{n}^{\prime},h_{n-1})}=\frac{\frac{1}{2}e^{\varepsilon(s_{n}^{\prime}-t(l_{n}))}}{1-\frac{1}{2}e^{-\varepsilon(s_{n}-t(l_{n}))}}\leq e^{\varepsilon d_{\mathbb{R}}(s_{n},s_{n}^{\prime})}.$
If $x_{n}=-1$, we similarly have
$\frac{\mathbb{P}(x_{n}=-1\mid s_{n},h_{n-1})}{\mathbb{P}(x_{n}=-1\mid s_{n}^{\prime},h_{n-1})}\leq e^{\varepsilon(s_{n}-s_{n}^{\prime})}.$
This completes the derivation, showing that the output distribution of the mechanism satisfies the upper bound
$\frac{\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})}{\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n})}\leq e^{\varepsilon d_{\mathbb{R}}(s_{n},s_{n}^{\prime})}$
in all possible cases.
We have shown that the smooth randomized response strategy satisfies the requirements of metric differential privacy. We now turn to the question of its optimality. In this context, optimality implies that the flipping probability should be minimized while still satisfying the metric differential privacy constraints.
Under metric differential privacy, if the distance between two signals $s_{n}$ and $s_{n}^{\prime}$ is close to zero, the probability of producing the same action under both signals should also be nearly the same. This condition imposes a critical constraint near the threshold $t(l_{n})$, where small changes in the signal can lead to different actions. Consequently, at the threshold $s_{n}=t(l_{n})$, the flipping probability must be exactly $\frac{1}{2}$ to guarantee the metric differential privacy.
For signals satisfying $s_{n}>t(l_{n})$ and $s_{n}^{\prime}>t(l_{n})$ (or $s_{n}<t(l_{n})$ and $s_{n}^{\prime}<t(l_{n})$) , to preserve metric differential privacy, the flipping probability function $u(s_{n})$ must decay no faster than $e^{-\varepsilon d(s_{n},t(l_{n}))}$. If $u(s_{n})$ decays more rapidly, the mechanism would violate the privacy constraint by allowing the output distribution to change too sharply. On the other hand, if $u(s_{n})$ decays more slowly, it introduces excessive randomness into the agent’s action, thereby reducing decision utility.
Therefore, the smooth randomized response strategy achieves an optimal balance: it provides just enough noise to satisfy the privacy constraints without injecting unnecessary randomness. This makes it an optimal choice for rational agents operating under metric differential privacy.
11 Proof of Theorem 4.8: Smoothly Asymptotic Learning
Without loss of generality, we assume $\theta=-1$. Then, the likelihood ratio of the public belief is defined as
$\pi_{n}:=\frac{\mathbb{P}(\theta=+1|x_{1},x_{2},\ldots,x_{n-1})}{\mathbb{P}(\theta=-1|x_{1},x_{2},\ldots,x_{n-1})}.$
which forms a martingale conditional on $\theta=-1$. To see this, we have
$\pi_{n+1}$ $=\frac{\mathbb{P}(\theta=+1|x_{1},x_{2},\ldots,x_{n-1})}{\mathbb{P}(\theta=-1|x_{1},x_{2},\ldots,x_{n-1})}$
$=\frac{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}|\theta=+1)}{\mathbb{P}(x_{1},x_{2},\ldots,x_{n-1}|\theta=-1)}$
$=\pi_{n}\frac{\mathbb{P}(x_{n}|\pi_{n},\theta=+1)}{\mathbb{P}(x_{n}|\pi_{n},\theta=-1)}.$
Therefore, we can compute the expectation as follows:
$E(\pi_{n+1}|\pi_{n},\theta=-1)=\pi_{n}\sum_{x_{n}\in\{0,1\}}\frac{\mathbb{P}(x_{n}|\pi_{n},\theta=+1)}{\mathbb{P}(x_{n}|\pi_{n},\theta=-1)}\mathbb{P}(x_{n}|\pi_{n},\theta=-1)=\pi_{n}.$
Since $\pi_{n}$ is a non-negative martingale and $\sup_{n}\mathbb{E}[\pi_{n}|\theta=-1]<\infty$, the martingale convergence theorem implies that, with probability one, the limit $\pi_{\infty}=\lim_{n\to\infty}\pi_{n}$ exists and is finite. Then a straightforward calculation shows that, when $x_{n}=+1$,
$l_{n+1}=l_{n}+\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}.$
Since the limit $l_{\infty}$ also exists and is finite, it follows that $\mathbb{P}(x_{n}=+1|l_{\infty},\theta=+1)=\mathbb{P}(x_{n}=+1|l_{\infty},\theta=-1)$ or $\mathbb{P}(x_{n}=-1|l_{\infty},\theta=+1)=\mathbb{P}(x_{n}=-1|l_{\infty},\theta=-1)$.
Assuming the private signals are Gaussian, when $\theta=+1$, the signals follow a normal distribution with mean +1+1 and variance $\sigma^{2}$, and when $\theta=-1$, they follow a normal distribution with mean 1-1 and variance $\sigma^{2}$. Thus,
$L_{n}=\log\frac{\exp\left(-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}\right)}{\exp\left(-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}\right)}=\frac{2s_{n}}{\sigma^{2}},$
which means $L_{t}$ is proportional to the signal $s_{t}$ and also normally distributed, conditioned on $\theta$, with variance $\frac{4}{\sigma^{2}}$. Therefore, if $s_{n}>-\sigma^{2}l_{n}/2=t(l_{n})$, the agent will choose $x_{n}=+1$; otherwise, the agent will choose $x_{n}=-1$.
To calculate $\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)$, we have
$\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)$ $=\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}$
$\quad+\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}.$
where $a=\frac{1}{2}$.
Similarly, for $\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$, we have
$\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ $=\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}$
$\quad+\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}.$
To compare the probabilities under different values of $\theta$, we define the following Gaussian densities:
$\phi_{+}(s_{n})=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}\right),\quad\phi_{-}(s_{n})=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}\right),$
which represent the conditional densities of $s_{n}$ given $\theta=+1$ and $\theta=-1$, respectively.
Let $g(s_{n})$ denote the following piecewise function:
$g(s_{n})=\begin{cases}1-ae^{\varepsilon(s_{n}+\frac{1}{2}\sigma^{2}l_{n})},&s_{n}<-\frac{1}{2}\sigma^{2}l_{n},\\ ae^{-\varepsilon(s_{n}+\frac{1}{2}\sigma^{2}l_{n})},&s_{n}>-\frac{1}{2}\sigma^{2}l_{n},\end{cases}$
Then, the conditional probabilities can be expressed as:
$\mathbb{P}(x_{n}=-1\mid l_{n},\theta=+1)=\int_{-\infty}^{\infty}\phi_{+}(s_{n})\cdot g(s_{n})\,ds_{n},\quad\mathbb{P}(x_{n}=-1\mid l_{n},\theta=-1)=\int_{-\infty}^{\infty}\phi_{-}(s_{n})\cdot g(s_{n})\,ds_{n}.$
We note that $g(s_{n})$ is a non-increasing function:
$s_{n}^{\prime}>s_{n}\quad\Rightarrow\quad g(s_{n}^{\prime})\leq g(s_{n}),$
and due to the exponential terms in its definition, there exists at least one pair $s_{n}^{\prime}>s_{n}$ such that
$g(s_{n}^{\prime})<g(s_{n}),$
i.e., $g(s_{n})$ is not constant almost everywhere.
Since $\phi_{+}$ and $\phi_{-}$ are Gaussian densities with the same variance but different means (+1+1 and 1-1, respectively), and $g(s_{n})$ is non-increasing and strictly decreasing on a set of nonzero measure, then for $l_{n}\in(-\infty,+\infty)$ and $\varepsilon>0$, we find that
$\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)-\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)<0.$ (15)
Thus, $\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)=\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ holds only if $l_{\infty}=$-$\infty$, since the limit $\pi_{\infty}$ exists and is finite.
Then, we find
$\lim_{n\to\infty}\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ $=\lim_{n\to\infty}\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}$
$\quad+\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}.$
$=\int_{-\infty}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ds_{n}-ae^{\varepsilon\sigma^{2}l_{\infty}/2}\int_{-\infty}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}e^{\varepsilon s_{n}}ds_{n}$
$=1-ae^{\varepsilon\sigma^{2}l_{\infty}/2}e^{\frac{\varepsilon^{2}\sigma^{2}}{2}}=1.$
12 Proof of Theorem 4.9: Learning Rate under Smooth Randomized Response
Before providing a formal proof of this result, we introduce a series of lemmas that will serve as foundational support for the main theorem.
Lemma 12.1
Let $X$ be a standard normal random variable, $\mathcal{N}(0,1)$. For $x>0$, the tail probability $\mathbb{P}(X\geq x)$ is bounded as follows:
$\frac{e^{-x^{2}/2}}{\sqrt{2\pi}}\left(\frac{1}{x}-\frac{1}{x^{3}}\right)\leq\mathbb{P}(X\geq x)\leq\frac{e^{-x^{2}/2}}{x\sqrt{2\pi}}.$
Proof.
We proceed by using integration by parts. First, we have
$\int_{x}^{\infty}e^{-\frac{t^{2}}{2}}dt=\left.\frac{-e^{-t^{2}/2}}{t\sqrt{2\pi}}\right|_{x}^{\infty}-\int_{x}^{\infty}\frac{e^{-t^{2}/2}}{t^{2}\sqrt{2\pi}}dt.$
Since the second term on the right-hand side is non-positive, we obtain the upper bound
$\int_{x}^{\infty}e^{-\frac{t^{2}}{2}}\,dt\leq\frac{e^{-x^{2}/2}}{x}.$
Thus,
$\mathbb{P}(X\geq x)\leq\frac{e^{-x^{2}/2}}{x\sqrt{2\pi}}.$
Next, we apply integration by parts to refine the lower bound:
$\int_{x}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-t^{2}/2}\,dt=\left.\frac{-e^{-t^{2}/2}}{t\sqrt{2\pi}}\right|_{x}^{\infty}+\left.\frac{e^{-t^{2}/2}}{t^{3}\sqrt{2\pi}}\right|_{x}^{\infty}+\int_{x}^{\infty}\frac{e^{-t^{2}/2}}{t^{4}\sqrt{2\pi}}\,dt.$
Dropping the last integral, which is non-negative, we obtain
$\mathbb{P}(X\geq x)\geq\frac{e^{-x^{2}/2}}{\sqrt{2\pi}}\left(\frac{1}{x}-\frac{1}{x^{3}}\right).$
Combining these bounds, we conclude that
$\frac{e^{-x^{2}/2}}{\sqrt{2\pi}}\left(\frac{1}{x}-\frac{1}{x^{3}}\right)\leq\mathbb{P}(X\geq x)\leq\frac{e^{-x^{2}/2}}{x\sqrt{2\pi}}.$
Lemma 12.2
For $\varepsilon\in(0,\infty)$,
$\lim_{l_{n}\to\infty}\frac{D_{+}(l_{n})}{G_{-}(-l_{n})}=1,$
where $D_{+}(l_{n})=\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}$, $G_{-}(-l_{n})=a(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}})e^{-\varepsilon\frac{\sigma^{2}l_{n}}{2}}$, and $a=\frac{1}{2}$.
Proof.
By definition, we have
$D_{+}(l_{n})=\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}=\log\frac{1-\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)}{1-\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)}.$
Using the approximation $\log(1-x)=-x+O(x^{2})$ for small $x$, we can rewrite $D_{+}(l_{n})$ as $D_{+}(l_{n})=\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)-\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)+O(\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)^{2})+O(\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)^{2})$.
We now analyze $\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ as follows:
$\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ $=\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}$
$\quad+\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}.$
Applying Lemma 12.1 and $1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2+1)}\geq\frac{1}{2}$, we establish both lower and upper bounds for the first term:
$\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}.$
We first present a lower bound:
$\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}\geq\frac{e^{-(-\frac{\sigma l_{n}}{2}+\frac{1}{\sigma})^{2}/2}}{\sqrt{2\pi}}\left(\frac{1}{-\frac{\sigma l_{n}}{2}+\frac{1}{\sigma}}-\frac{1}{(-\frac{\sigma l_{n}}{2}+\frac{1}{\sigma})^{3}}\right).$
Next, we provide an upper bound:
$\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}\leq\frac{e^{-(-\frac{\sigma l_{n}}{2}+\frac{1}{\sigma})^{2}/2}}{(-\frac{\sigma l_{n}}{2}+\frac{1}{\sigma})\sqrt{2\pi}}.$
Thus, as $l_{n}\to\infty$, the first term is $\Theta\left(\frac{e^{-\frac{\sigma^{2}l_{n}^{2}}{4}}}{l_{n}}\right)$.
For the second term, we have:
$\quad\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}$
$=ae^{-\varepsilon\sigma^{2}l_{n}/2}\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon s_{n}}ds_{n}$
$\xrightarrow{l_{n}\to\infty}ae^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}e^{-\varepsilon\frac{\sigma^{2}l_{n}}{2}}.$
Similarly, for $\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)$, we have
$\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ $=\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}$
$\quad+\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}.$
For the first term, we establish both lower and upper bounds for the integral:
$\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}.$
We first present the lower bound:
$\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}\geq\frac{e^{-(-\frac{\sigma l_{n}}{2}-\frac{1}{\sigma})^{2}/2}}{\sqrt{2\pi}}\left(\frac{1}{-\frac{\sigma l_{n}}{2}-\frac{1}{\sigma}}-\frac{1}{(-\frac{\sigma l_{n}}{2}-\frac{1}{\sigma})^{3}}\right).$
Next, we provide the corresponding upper bound:
$\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}\leq\frac{e^{-(-\frac{\sigma l_{n}}{2}-\frac{1}{\sigma})^{2}/2}}{(-\frac{\sigma l_{n}}{2}-\frac{1}{\sigma})\sqrt{2\pi}}.$
Thus, as $l_{n}\to\infty$, the first term is $\Theta\left(\frac{e^{-\frac{\sigma^{2}l_{n}^{2}}{4}}}{l_{n}}\right)$.
For the second term, we have:
$\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}\xrightarrow{l_{n}\to\infty}ae^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}e^{-\varepsilon\frac{\sigma^{2}l_{n}}{2}}.$
Thus, as $l_{n}\to\infty$, both $\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ and $\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)$ are dominated by the second term.
Therefore, we have
$\lim_{l_{n}\to\infty}\frac{D_{+}(l_{n})}{G_{-}(-l_{n})}=1,$
where $D_{+}(l_{n})=\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}$, $G_{-}(-l_{n})=\frac{1}{2}(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}})e^{-\varepsilon\frac{\sigma^{2}l_{n}}{2}}$. ∎
Lemma 12.3
Let $A,B:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ be continuous functions that are eventually monotonically decreasing and tend to zero. Consider the sequences $(a_{t})$ and $(b_{t})$ defined by the recurrence relations:
$a_{t+1}=a_{t}+A(a_{t})\quad\text{and}\quad b_{t+1}=b_{t}+B(b_{t}).$
Suppose that
$\lim_{x\to\infty}\frac{A(x)}{B(x)}=1.$
Then it follows that
$\lim_{t\to\infty}\frac{a_{t}}{b_{t}}=1.$
This is proved by Hann-Caruthers et al. (2018, Lemma 9).
Lemma 12.4
Assume that $A:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ is a continuous function with a convex differentiable tail and that $A(x)$ approaches zero as $x$ tends to infinity. Let $(a_{n})$ be a sequence satisfying the recurrence relation:
$a_{n+1}=a_{n}+A(a_{n}).$
Moreover, suppose there exists a function $f:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ such that $f^{\prime}(t)=A(f(t))$ for sufficiently large $t$. Then we have
$\lim_{n\to\infty}\frac{f(n)}{a_{n}}=1.$
This proved by Hann-Caruthers et al. (2018, Lemma 10).
Given these lemmas, we are now prepared to proceed with the proof of Theorem 4.9.
We begin by solving the differential equation
$\frac{\mathrm{d}f}{\mathrm{d}t}(t)=G_{-}(-f(t)),$
where $G_{-}(-f(t))=Ce^{-\varepsilon\frac{\sigma^{2}f(t)}{2}}$, and $C=\frac{1}{2}(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}})$. We proceed by separating variables:
$\int e^{\varepsilon\frac{\sigma^{2}f(t)}{2}}\,df=C\int\,dt.$
Integrating both sides, we obtain the following:
$\frac{2}{\varepsilon\sigma^{2}}e^{\varepsilon\frac{\sigma^{2}f(t)}{2}}=Ct,$
Solving for $f(t)$, we find:
$f(t)=\frac{2}{\varepsilon\sigma^{2}}\log(\frac{\varepsilon\sigma^{2}}{2}Ct).$
Let $(a_{n})$ be a sequence in $\mathbb{R}_{>0}$ defined by the recurrence relation:
$a_{n+1}=a_{n}+G(-a_{n}).$
According to Lemma 12.4, this sequence $(a_{n})$ is closely approximated by the function $f(n)$, which satisfies the associated differential equation, and we have
$\lim_{n\to\infty}\frac{a_{n}}{f(n)}=1.$
Now, assuming that $\theta=+1$, all agents eventually choose action +1+1 starting from some time onward, with probability 1. Therefore, with probability 1, the recurrence for $\ell_{n}$ can be expressed as
$\ell_{n+1}=\ell_{n}+D_{+}(\ell_{n}),$
for all sufficiently large $n$. Additionally, by Lemma 12.2, it follows that
$\lim_{x\to\infty}\frac{D_{+}(x)}{G(-x)}=1.$
Thus, applying Lemma 12.3, we conclude that
$\lim_{n\to\infty}\frac{\ell_{n}}{a_{n}}=1$
with probability 1. Consequently, we can combine the limits to obtain
$\lim_{n\to\infty}\frac{\ell_{n}}{f(n)}=\lim_{n\to\infty}\frac{\ell_{n}}{a_{n}}\cdot\frac{a_{n}}{f(n)}=1,$
with probability 1, as desired.
13 Proof of Theorem 4.10: Finite Expected Time to First Correct Action
Before presenting the formal proof of Theorem 4.11 and Theorem 4.11, we first establish several lemmas that will be used in the proof.
Lemma 13.1
For each n, $\pi_{n}\geq\frac{1}{2}$ if $x_{n}=+1$ and $\pi_{n}\leq\frac{1}{2}$ if $x_{n}=-1$.
Proof.
By Bayes’ rule, we express the ratio of posterior probabilities as follows:
$\frac{\pi_{n}}{1-\pi_{n}}$ $=\frac{\mathbb{P}(\theta=+1|x_{1},\dots,x_{n})}{\mathbb{P}(\theta=-1|x_{1},\dots,x_{n})}$
$=\frac{\mathbb{P}(x_{1},\dots,x_{n}|\theta=+1)}{\mathbb{P}(x_{1},\dots,x_{n}|\theta=-1)}$
$=\frac{\mathbb{P}(x_{n}|\theta=+1,x_{1},\dots,x_{n-1})\mathbb{P}(x_{1},\dots,x_{n-1}|\theta=+1)}{\mathbb{P}(x_{n}|\theta=-1,x_{1},\dots,x_{n-1})\mathbb{P}(x_{1},\dots,x_{n-1}|\theta=-1)}.$
Considering the smooth randomized response strategy, we further expand:
$\frac{\pi_{n}}{1-\pi_{n}}$ $=\frac{(1-u_{n})\mathbb{P}(a_{n}|\theta=+1,x_{1},\dots,x_{n-1})+u_{n}\mathbb{P}(\tilde{a}_{n}|\theta=+1,x_{1},\dots,x_{n-1})}{(1-u_{n})\mathbb{P}(a_{n}|\theta=-1,x_{1},\dots,x_{n-1})+u_{n}\mathbb{P}(\tilde{a}_{n}|\theta=-1,x_{1},\dots,x_{n-1})}$
$=\frac{(1-2u_{n})\frac{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=+1)}{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=-1)}+\frac{u_{n}}{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=-1)}}{1-2u_{n}+\frac{u_{n}}{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=-1)}}.$
Next, we compute the conditional probability of $\theta=+1$ given the past observations and the current action before the randomized response:
$\mathbb{P}(\theta=+1|x_{1},\dots,x_{n-1},a_{n})$ $=\mathbb{E}[\mathbf{1}_{\theta=+1}|x_{1},\dots,x_{n-1},a_{n}]$
$=\mathbb{E}\Big[\mathbb{E}[\mathbf{1}_{\theta=+1}|x_{1},\dots,x_{n-1},a_{n},s_{n}]\Big|x_{1},\dots,x_{n-1},a_{n}\Big]$
$=\mathbb{E}[\mathbb{P}(\theta=+1|x_{1},\dots,x_{n-1},a_{n},s_{n})|x_{1},\dots,x_{n}].$
where the second equality follows from the law of iterated conditional expectations. Since the decision rule dictates that $a_{n}=+1$ if and only if $\mathbb{P}(\theta=+1|x_{1},\dots,x_{n-1},a_{n},s_{n})\geq\frac{1}{2}$, it follows that: If $a_{n}=+1$, then $\mathbb{P}(\theta=+1|x_{1},\dots,x_{n-1},a_{n})\geq\frac{1}{2}$. If $a_{n}=-1$, then $\mathbb{P}(\theta=+1|x_{1},\dots,x_{n-1},a_{n})\leq\frac{1}{2}$.
Thus, when $a_{n}=+1$, we have:
$\frac{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=+1)}{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=-1)}=\frac{\mathbb{P}(\theta=+1|x_{1},\dots,x_{n-1},a_{n})}{\mathbb{P}(\theta=-1|x_{1},\dots,x_{n-1},a_{n})}\geq 1.$
Similarly, when $a_{n}=-1$, we obtain:
$\frac{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=+1)}{\mathbb{P}(x_{1},\dots,x_{n-1},a_{n}|\theta=-1)}=\frac{\mathbb{P}(\theta=+1|x_{1},\dots,x_{n-1},a_{n})}{\mathbb{P}(\theta=-1|x_{1},\dots,x_{n-1},a_{n})}\leq 1.$
Since $0\leq u_{n}\leq\frac{1}{2}$, we conclude that: If $a_{n}=+1$, then
$\frac{\pi_{n}}{1-\pi_{n}}\geq 1.$
If $a_{n}=-1$, then
$\frac{\pi_{n}}{1-\pi_{n}}\leq 1.$
Thus, we conclude that $\pi_{n}\geq\frac{1}{2}$ when $x_{n}=+1$ and $\pi_{n}\leq\frac{1}{2}$ when $x_{n}=-1$, completing the proof. ∎
Lemma 13.2
For any $\pi\in(0,1)$, the expected stopping time satisfies
$\mathbb{E}_{\pi,-}[\tau]=\frac{\pi}{1-\pi}\mathbb{E}[\tau].$
Proof.
Since the states $\theta=+1$ and $\theta=-1$ are ex ante equally likely, the posterior odds ratio satisfies:
$\frac{\pi^{*}_{n+1}}{1-\pi^{*}_{n+1}}=\frac{\mathbb{P}_{+}(x_{1}=\dots=x_{n}=+1)}{\mathbb{P}_{-}(x_{1}=\dots=x_{n}=+1)}.$
From this, we have:
$u_{n}=\mathbb{P}_{-}(\tau>n)=\mathbb{P}_{-}(x_{1}=\dots=x_{n}=+1)=\frac{1-\pi^{*}_{n+1}}{\pi^{*}_{n+1}}\mathbb{P}_{+}(x_{1}=\dots=x_{n}=+1).$
For the general case where $\pi\neq\frac{1}{2}$, we extend the previous argument to obtain:
$\frac{\pi^{*}_{n+1}}{1-\pi^{*}_{n+1}}=\frac{\pi}{1-\pi}\frac{\mathbb{P}(x_{1}=\dots=x_{n}=+1|\theta=+1)}{\mathbb{P}(x_{1}=\dots=x_{n}=+1|\theta=-1)}.$
Similarly, defining:
$\tilde{u}_{n}=\mathbb{P}_{\pi,-}(\tau>n)=\frac{\pi}{1-\pi}\frac{1-\pi^{*}_{n+1}}{\pi^{*}_{n+1}}\mathbb{P}_{+}(x_{1}=\dots=x_{n}=+1),$
we derive the expectation:
$\mathbb{E}_{\pi,-}[\tau]=\sum_{n=1}^{\infty}\tilde{u}_{n}=\frac{\pi}{1-\pi}\sum_{n=1}^{\infty}u_{n}=\frac{\pi}{1-\pi}\mathbb{E}_{\pi}[\tau].$
This completes the proof. ∎
Lemma 13.3
There exists a constant $a>0$ such that
$\mathbb{P}_{\pi,-}(x_{n}=-1,\text{ for all }n)\geq a,$
for all $\pi\leq\frac{1}{2}$.
Proof.
Define the stopping time of the first incorrect choice as
$\sigma:=\inf\{n:x_{n}\neq\theta\}.$
Let $Y_{n}:=\frac{\pi_{n}}{1-\pi_{n}}$. Since $Y_{n}$ is a positive martingale under $\mathbb{P}_{\pi,-}$, we apply the optional stopping theorem at $\sigma$, obtaining:
$\mathbb{E}_{\pi,-}[Y_{\sigma}]\leq Y_{1}.$
Thus, we have:
$\mathbb{E}_{\pi,-}[Y_{\sigma}\mathbf{1}_{\sigma<+\infty}]\leq Y_{1}.$
Since $x_{\sigma-1}=+1$, Lemma 13.1 implies that $Y_{\sigma}\geq 1$ almost surely. This result yields:
$\mathbb{P}_{\pi,-}(\sigma<+\infty)\leq Y_{1}=\frac{\pi}{1-\pi}.$
Moreover, for $\pi\leq\frac{1}{2}-\alpha$, we have the lower bound:
$\mathbb{P}_{\pi,-}(\sigma=+\infty)\geq\frac{4\alpha}{1+2\alpha}.$
For $\pi\in[\frac{1}{2}-\alpha,\frac{1}{2}]$, there exists a constant $m>0$ such that
$\mathbb{P}_{\pi,-}(x_{1}=-1)\geq m.$
Define $\phi(\pi|x)$ as the public belief of the next agent when the current agent has belief $\pi$ and takes action $x$. By Bayes’ rule, we obtain:
$\frac{\phi(\pi|x)}{1-\phi(\pi|x)}=\frac{\pi}{1-\pi}\cdot\frac{\mathbb{P}(x_{n}=-1|\pi,\theta=+1)}{\mathbb{P}(x_{n}=-1|\pi,\theta=-1)}.$
From Equation 15, we know that
$\mathbb{P}(x_{n}=-1|\pi,\theta=+1)<\mathbb{P}(x_{n}=-1|\pi,\theta=-1),$
which implies that $\phi(\pi|x)<\pi$. Since $\pi\in[\frac{1}{2}-\alpha,\frac{1}{2}]$, it follows that
$\phi(\pi|x)\leq\frac{1}{2}-\alpha.$
By the Markov property of $\pi_{n}$, we deduce:
$\mathbb{P}_{\pi,-}(\sigma=+\infty)=\mathbb{P}_{\pi,-}(x_{1}=-1)\cdot\mathbb{P}_{\phi(\pi|x),-}(\sigma=+\infty)\geq m\cdot\frac{4\alpha}{1+2\alpha}.$
Thus, there exists a constant $a>0$ such that
$\mathbb{P}_{\pi,-}(x_{n}=-1,\text{ for all }n)\geq a,$
for all $\pi\leq\frac{1}{2}$, completing the proof. ∎
To complete the proof of Theorem 4.10, it remains to establish Equation 10. To proceed, recall the definition of $u_{n}$:
$u_{n}=\prod_{k=1}^{n}\mathbb{P}(x_{k}=+1|l_{k},\theta=-1).$
By Equation 8, we can express $u_{n}$ as:
$u_{n}=\prod_{k=2}^{n}\mathbb{P}(x_{k}=+1|l_{k},\theta=-1)\frac{1-\pi_{2}^{*}}{\pi_{2}^{*}}\mathbb{P}(x_{1}=+1|l_{1},\theta=+1),$
and further as:
$u_{n}=\prod_{k=3}^{n}\mathbb{P}(x_{k}=+1|l_{k},\theta=-1)\frac{1-\pi_{3}^{*}}{\pi_{3}^{*}}\mathbb{P}(x_{2}=+1|l_{2},\theta=+1)\mathbb{P}(x_{1}=+1|l_{1},\theta=+1).$
Then
$u_{n}=\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)\frac{1-\pi_{n}^{*}}{\pi_{n}^{*}}\prod_{k=1}^{n-1}\mathbb{P}(x_{k}=+1|l_{k},\theta=+1).$
Using the recurrence relationship for $\pi_{n+1}^{*}$, this can be further simplified as:
$u_{n}=\frac{1-\pi_{n+1}^{*}}{\pi_{n+1}^{*}}\prod_{k=1}^{n}\mathbb{P}(x_{k}=+1|l_{k},\theta=+1)$
By Lemma 13.1, $\pi_{n+1}^{*}\geq\frac{1}{2}$, and by Lemma 13.3, we have
$c(1-\pi_{n+1}^{*})\leq u_{n}\leq 2(1-\pi_{n+1}^{*}).$
Since $l_{n}=\log\left(\frac{\pi_{n}^{*}}{1-\pi_{n}^{*}}\right)$, we have $1-\pi_{n}^{*}\sim e^{-l_{n}}$, where $l_{n}=\frac{2}{\varepsilon\sigma^{2}}\log(C(\varepsilon)n)+c$, where $c$ is a constant and
$C(\varepsilon)=\frac{\varepsilon^{2}\sigma^{2}}{4}\left(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}\right)$
Thus, the expected time $\mathbb{E}[\tau]$ until the first correct action satisfies
$\mathbb{E}[\tau]=\sum_{n=1}^{\infty}u_{n}=C_{1}\left(C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\zeta\left(\frac{2}{\varepsilon\sigma^{2}}\right)\right),$
where $C_{1}$ is a positive constant and $\zeta\left(x\right)=\sum_{n=1}^{\infty}n^{-x}$, for some constant $C(\varepsilon)$ depending on $\varepsilon$. Since $\frac{2}{\varepsilon\sigma^{2}}>1$, the series converges, and thus the expected time until the first correct action is finite:
$\mathbb{E}[\tau]<+\infty.$
14 Proof of Theorem 4.11: Finite Expected Total Number of Incorrect Actions
Based on the above lemmas and Theorem 4.10, we now present the formal proof of Theorem 4.11. Without loss of generality, assume $\theta=-1$. We define $\sigma_{1}=1$ and let $\tau_{1}=\inf\{n:x_{n}=-1\}$ be the starting index of the first incorrect and correct runs. For $k>1$, define recursively:
$\sigma_{k}:=\inf\{n>\tau_{k-1}:x_{n}=+1\},\quad\tau_{k}:=\inf\{n>\sigma_{k}:x_{n}=-1\}.$
Then, we define the length of the $k$th bad run as $\Delta_{k}:=\tau_{k}-\sigma_{k}$. By monotone convergence theorem,
$\mathbb{E}_{-}[W]=\sum_{k=1}^{+\infty}\mathbb{E}_{-}\big[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}\big].$
The indicator function $\mathbf{1}_{\sigma_{k}<+\infty}$ ensures that only finite stopping times $\sigma_{k}$ contribute to the expectation. This accounts for the possibility that, after some point, agents may always take the correct action, leading to $\sigma_{k}=+\infty$ and thus excluding those terms from the summation.
For $k\geq 2$, we derive an upper bound for $\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}|\mathcal{H}_{\tau_{k-1}}]$, where $\mathcal{H}_{\tau_{k-1}}$ is the $\sigma$-algebra at time $\tau_{k-1}$.
$\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}|\mathcal{H}_{\tau_{k-1}}]$ $=\mathbb{E}_{-}\Big[\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}|\mathcal{H}_{\sigma_{k}}]\Big|\mathcal{H}_{\tau_{k-1}}\Big]$
$=\mathbb{E}_{-}\Big[\mathbf{1}_{\sigma_{k}<+\infty}\mathbb{E}_{-}[\Delta_{k}|\mathcal{H}_{\sigma_{k}}]\Big|\mathcal{H}_{\tau_{k-1}}\Big]$
$=\mathbb{E}_{-}\Big[\mathbf{1}_{\sigma_{k}<+\infty}\mathbb{E}_{\pi_{\sigma_{k}},-}[\tau]\Big|\mathcal{H}_{\tau_{k-1}}\Big]$
$\leq\mathbb{E}[\tau]\mathbb{E}_{-}\Big[\frac{\pi_{\sigma_{k}}}{1-\pi_{\sigma_{k}}}\mathbf{1}_{\sigma_{k}<+\infty}\Big|\mathcal{H}_{\tau_{k-1}}\Big]$
$\leq\mathbb{E}[\tau]\frac{\pi_{\tau_{k-1}}}{1-\pi_{\tau_{k-1}}}$
$\leq\mathbb{E}[\tau].$
where the first equality follows from the law of iterated expectations, the second equality holds since $\sigma_{k}$ is $\mathcal{H}_{\sigma_{k}}$-measurable, the third equality follows from the Markov property of $(\pi_{n})$, the inequality follows from Lemma 13.2 , the next step is derived using the Doob optional sampling theorem for positive martingales, and the final inequality holds by Lemma 13.1.
Then we derive an upper bound for $\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}]$ as follows:
$\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}]$ $=\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}\mathbf{1}_{\tau_{k-1}<+\infty}]$
$=\mathbb{E}_{-}\Big[\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}\mathbf{1}_{\tau_{k-1}<+\infty}\Big|\mathcal{H}_{\tau_{k-1}}]\Big]$
$=\mathbb{E}_{-}\Big[\mathbf{1}_{\tau_{k-1}<+\infty}\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}\Big|\mathcal{H}_{\tau_{k-1}}]\Big]$
$\leq\mathbb{E}[\tau]\mathbb{P}_{-}(\tau_{k-1}<+\infty)$
$\leq\mathbb{E}[\tau]\mathbb{P}_{-}(\sigma_{k-1}<+\infty).$
where the first equality holds since $\sigma_{k}<+\infty$ implies $\tau_{k-1}<+\infty$, the second equality follows from the law of iterated expectations, the third equality holds since $\tau_{k-1}$ is $\mathcal{H}_{\tau_{k-1}}$ -measurable. Then we sum $\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}]$ over $k\geq 1$,
$\mathbb{E}_{-}[W]\leq\mathbb{E}[\tau]+\mathbb{E}[\tau]\sum_{k=1}^{+\infty}{P}_{-}(\sigma_{k}<+\infty).$
Then, we derive the inequality recurrence relationship for $\mathbb{P}_{-}(\sigma_{k}<+\infty)$:
$\mathbb{P}_{-}(\sigma_{k}<+\infty)$ $=\mathbb{E}_{-}\Big[\mathbb{E}_{-}[\mathbf{1}_{\sigma_{k}<+\infty}|\mathcal{H}_{\tau_{k-1}}]\Big]$
$=\mathbb{E}_{-}\Big[\big(1-\mathbb{P}_{\pi_{\tau_{k}-1},L}(x_{n}=l,\text{ for all }n)\big)\mathbf{1}_{\tau_{k-1}<+\infty}\Big]$
$\leq(1-a)\mathbb{P}_{-}(\tau_{k-1}<+\infty)$
$\leq(1-a)\mathbb{P}_{-}(\sigma_{k-1}<+\infty).$
where the first inequality follows from Lemma 13.3. Since $\mathbb{P}_{-}(\sigma_{k}<+\infty)\leq(1-c)^{k-1}$, then $\mathbb{E}[\tau]\leq\mathbb{E}_{-}[W]\leq C_{3}\mathbb{E}[\tau]$.
15 The Evolution of Log-Likelihood Ratio
Figure 15.1: (a) Evolution of $l_{n}$ over steps $n$ for different privacy budgets and non-private scenario $\varepsilon$. (b) Comparison between exact and approximate solutions for $l_{n}$ when $\varepsilon=1$.
To support Theorem 4.9, we performed simulations to analyze the evolution of the log-likelihood ratio $l_{n}$ without an approximation. Figure 15.1(a) illustrates the trajectory of $l_{n}$ under three different privacy budgets and the non-private scenario, revealing a clear relationship between privacy constraints and learning efficiency. Specifically, as the privacy budget decreases, the rate at which the log-likelihood ratio increases becomes steeper. This observation suggests that stronger privacy constraints, while introducing noise into individual actions, paradoxically enhance the speed of belief convergence by amplifying the informativeness of correct actions. This aligns with our theoretical findings that the smooth randomized response mechanism asymmetrically affects the likelihood of correct and incorrect actions, thus improving the efficiency of information aggregation. Moreover, we observe a significant difference in learning speed between the private and non-private settings: In the private setting, the log-likelihood ratio grows at a rate of $\Theta_{\varepsilon}(\log(n))$, whereas in the non-private setting, where privacy budgets approach infinity, asymptotic learning occurs at the slower rate of $\Theta(\sqrt{\log(n)})$. This confirms that our privacy-aware mechanism leads to a provable acceleration of information aggregation compared to classical sequential learning models.
Furthermore, we validate the robustness of our approximation approach by comparing the exact evolution of $l_{n}$ with its approximation, as depicted in Figure 15.1(b). The results indicate that, while there are minor deviations, overall trends remain consistent across different privacy settings. This confirms that the approximation method used in our theoretical analysis reliably captures the key dynamics of sequential learning under privacy constraints. These findings further reinforce the counterintuitive insight that privacy-preserving mechanisms, rather than simply introducing noise that degrades learning performance, can strategically reshape the information structure to accelerate the convergence of beliefs in sequential decision-making settings.
16 Proof of Theorem 4.12: Asymptotic Learning under Heterogeneous Privacy Budgets
Based on a similar argument to the proof of asymptotic learning in the homogeneous setting, we assume without loss of generality that the true state is $\theta=-1$. In this case, we can show that $\pi_{n}$ is a non-negative martingale under the true state and satisfies $\sup_{n}\mathbb{E}[\pi_{n}\mid\theta=-1]<\infty$. By the martingale convergence theorem, it follows that the limit $\pi_{\infty}=\lim_{n\to\infty}\pi_{n}$ exists almost surely and is finite.
Moreover, from the update rule of the public log-likelihood ratio:
$l_{n+1}=l_{n}+\log\frac{\mathbb{P}(x_{n}=+1\mid l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1\mid l_{n},\theta=-1)},$
the convergence of $\pi_{n}$ implies that $l_{n}$ also converges almost surely to a finite limit $l_{\infty}$. Thus, in the limit we must have
$\mathbb{P}(x_{n}=+1\mid l_{\infty},\theta=+1)=\mathbb{P}(x_{n}=+1\mid l_{\infty},\theta=-1),$
or equivalently,
$\mathbb{P}(x_{n}=-1\mid l_{\infty},\theta=+1)=\mathbb{P}(x_{n}=-1\mid l_{\infty},\theta=-1).$
To compute $\mathbb{P}(x_{n}=-1\mid l_{n},\theta=+1)$, we have:
$\mathbb{P}(x_{n}=-1\mid l_{n},\theta=+1)$ $=\int_{0}^{1}\int_{-\infty}^{-\frac{\sigma^{2}l_{n}}{2}}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon_{n}(s_{n}+\frac{\sigma^{2}l_{n}}{2})}\right)ds_{n}d\varepsilon_{n}$
$\quad+\int_{0}^{1}\int_{-\frac{\sigma^{2}l_{n}}{2}}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon_{n}(s_{n}+\frac{\sigma^{2}l_{n}}{2})}ds_{n}d\varepsilon_{n}$
$=\int_{-\infty}^{-\frac{\sigma^{2}l_{n}}{2}}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\int_{0}^{1}\left(1-ae^{\varepsilon_{n}(s_{n}+\frac{\sigma^{2}l_{n}}{2})}\right)d\varepsilon_{n}ds_{n}$
$\quad+\int_{-\frac{\sigma^{2}l_{n}}{2}}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\int_{0}^{1}ae^{-\varepsilon_{n}(s_{n}+\frac{\sigma^{2}l_{n}}{2})}d\varepsilon_{n}ds_{n}.$
Similarly, for $\mathbb{P}(x_{n}=-1\mid l_{n},\theta=-1)$:
$\mathbb{P}(x_{n}=-1\mid l_{n},\theta=-1)$ $=\int_{-\infty}^{-\frac{\sigma^{2}l_{n}}{2}}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\int_{0}^{1}\left(1-ae^{\varepsilon_{n}(s_{n}+\frac{\sigma^{2}l_{n}}{2})}\right)d\varepsilon_{n}ds_{n}$
$\quad+\int_{-\frac{\sigma^{2}l_{n}}{2}}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\int_{0}^{1}ae^{-\varepsilon_{n}(s_{n}+\frac{\sigma^{2}l_{n}}{2})}d\varepsilon_{n}ds_{n}.$
Define the signal densities:
$\phi_{+}(s_{n})=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}\right),\quad\phi_{-}(s_{n})=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}\right),$
and define the piecewise function
$g(s_{n})=\begin{cases}\mathbb{E}_{\varepsilon_{n}}[1-ae^{\varepsilon_{n}(s_{n}+\frac{1}{2}\sigma^{2}l_{n})}],&s_{n}<-\frac{1}{2}\sigma^{2}l_{n},\\ \mathbb{E}_{\varepsilon_{n}}[ae^{-\varepsilon_{n}(s_{n}+\frac{1}{2}\sigma^{2}l_{n})}],&s_{n}\geq-\frac{1}{2}\sigma^{2}l_{n}.\end{cases}$
Then we can write:
$\mathbb{P}(x_{n}=-1\mid l_{n},\theta=+1)=\int_{-\infty}^{\infty}\phi_{+}(s_{n})\cdot g(s_{n})\,ds_{n},\quad\mathbb{P}(x_{n}=-1\mid l_{n},\theta=-1)=\int_{-\infty}^{\infty}\phi_{-}(s_{n})\cdot g(s_{n})\,ds_{n}.$
Since $g(s_{n})$ is non-increasing and strictly decreasing on a set of positive measure, and $\phi_{+}$ and $\phi_{-}$ are Gaussians with the same variance but different means, it follows that
$\mathbb{P}(x_{n}=-1\mid l_{n},\theta=+1)<\mathbb{P}(x_{n}=-1\mid l_{n},\theta=-1).$ (16)
Therefore, the equality $\mathbb{P}(x_{n}=-1\mid l_{\infty},\theta=+1)=\mathbb{P}(x_{n}=-1\mid l_{\infty},\theta=-1)$ holds only if $l_{\infty}=$-$\infty$.
Finally, we compute the limiting behavior:
$\lim_{n\to\infty}\mathbb{P}(x_{n}=-1\mid l_{n},\theta=-1)$
$=\int_{0}^{1}\int_{-\infty}^{\infty}\phi_{-}(s_{n})\left[\mathbf{1}_{\{s_{n}<-\frac{1}{2}\sigma^{2}l_{n}\}}\left(1-ae^{\varepsilon_{n}(s_{n}+\frac{1}{2}\sigma^{2}l_{n})}\right)+\mathbf{1}_{\{s_{n}\geq-\frac{1}{2}\sigma^{2}l_{n}\}}ae^{-\varepsilon_{n}(s_{n}+\frac{1}{2}\sigma^{2}l_{n})}\right]ds_{n}d\varepsilon_{n}$
$=1-a\int_{0}^{1}e^{\varepsilon_{n}\sigma^{2}l_{\infty}/2}\int_{-\infty}^{\infty}\phi_{-}(s_{n})e^{\varepsilon_{n}s_{n}}ds_{n}d\varepsilon_{n}$
$\geq 1-ae^{\sigma^{2}l_{\infty}/2}\cdot\mathbb{E}_{\varepsilon_{n}}\left[e^{\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}\right]=1.$
17 Proof of Theorem 4.13: Learning Rate under Smooth Randomized Response with Heterogeneous Privacy Budget
Before providing a formal proof of this result, we introduce a series of lemmas that will serve as foundational support for the main theorem.
Proof.
By definition, we have
$D_{+}(l_{n})=\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}=\log\frac{1-\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)}{1-\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)}.$
Using the approximation $\log(1-x)=-x+O(x^{2})$ for small $x$, we can rewrite $D_{+}(l_{n})$ as
$D_{+}(l_{n})=\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)-\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)+O(\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)^{2})+O(\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)^{2}).$
We now analyze $\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ as follows:
$\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ $=\int_{0}^{1}\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon_{n}(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}d\varepsilon_{n}$
$\quad+\int_{0}^{1}\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon_{n}(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}d\varepsilon_{n}.$
Applying Lemma 12.1 and $1-ae^{\varepsilon(s_{n}+\sigma^{2}l_{n}/2+1)}\geq\frac{1}{2}$, we establish bounds for the first term by the same argument in Lemma 12.2. Thus, as $l_{n}\to\infty$, the first term is $\Theta\left(\frac{e^{-\frac{\sigma^{2}l_{n}^{2}}{4}}}{l_{n}}\right)$.
For the second term, we have:
$\quad\lim_{l_{n}\to\infty}\int_{0}^{1}\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon_{n}(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}d\varepsilon_{n}$
$=\lim_{l_{n}\to\infty}\int_{0}^{1}ae^{-\varepsilon_{n}\sigma^{2}l_{n}/2}\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}+1)^{2}}{2\sigma^{2}}}e^{-\varepsilon_{n}s_{n}}ds_{n}d\varepsilon_{n}$
$=\mathbb{E}_{\varepsilon_{n}}\left[ae^{\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right].$
The third equality follows from the Lebesgue dominated convergence theorem.
Similarly, for $\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)$, we have
$\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)$ $=\int_{0}^{1}\int_{-\infty}^{-\sigma^{2}l_{n}/2}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}\left(1-ae^{\varepsilon_{n}(s_{n}+\sigma^{2}l_{n}/2)}\right)ds_{n}d\varepsilon_{n}$
$\quad+\int_{0}^{1}\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon_{n}(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}d\varepsilon_{n}.$
As $l_{n}\to\infty$, the first term is again $\Theta\left(\frac{e^{-\frac{\sigma^{2}l_{n}^{2}}{4}}}{l_{n}}\right)$. For the second term, we have:
$\lim_{l_{n}\to\infty}\int_{0}^{1}\int_{-\sigma^{2}l_{n}/2}^{\infty}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(s_{n}-1)^{2}}{2\sigma^{2}}}ae^{-\varepsilon_{n}(s_{n}+\sigma^{2}l_{n}/2)}ds_{n}d\varepsilon_{n}=\mathbb{E}_{\varepsilon_{n}}\left[ae^{-\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right].$
Thus, as $l_{n}\to\infty$, both $\mathbb{P}(x_{n}=-1|l_{n},\theta=-1)$ and $\mathbb{P}(x_{n}=-1|l_{n},\theta=+1)$ are dominated by the second term. Therefore,
$\lim_{l_{n}\to\infty}\frac{D_{+}(l_{n})}{G_{-}(-l_{n})}=1,$
where
$D_{+}(l_{n})=\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)},\quad G_{-}(-l_{n})=\frac{1}{2}\mathbb{E}_{\varepsilon_{n}}\left[\left(e^{\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}-e^{-\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}\right)e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right].$
Since the support of $\varepsilon_{n}$ is $[0,1]$, and the event $\{\varepsilon_{n}=0\}$ has measure zero under the uniform distribution, it follows that
$\tilde{C}_{1}\,\mathbb{E}_{\varepsilon_{n}}\left[e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right]\leq G_{-}(-l_{n})\leq\tilde{C}_{2}\,\mathbb{E}_{\varepsilon_{n}}\left[e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right],$
where $\tilde{C}_{1},\tilde{C}_{2}>0$ are constants independent of $n$. Since $\mathbb{E}_{\varepsilon_{n}}[e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}]$ is the moment generating function of the uniform distribution, we have
$\mathbb{E}_{\varepsilon_{n}}\left[e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right]=\frac{2}{\sigma^{2}l_{n}}.$
Therefore,
$\lim_{l_{n}\to\infty}\frac{D_{+}(l_{n})}{G_{-}(-l_{n})}=1,$
where $G_{-}(-l_{n})=\frac{\tilde{C}}{\sigma^{2}l_{n}}$ for some constant $\tilde{C}$.
Given the above lemma, we are now prepared to proceed with the proof of Theorem 4.13.
We begin by solving the differential equation
$\frac{\mathrm{d}f}{\mathrm{d}t}(t)=G_{-}(-f(t)),$ (17)
where
$G_{-}(-f(t))=\frac{\tilde{C}}{\sigma^{2}f(t)}.$
We proceed by separating variables:
$\int\sigma^{2}f(t)\,df=\tilde{C}\int\,dt.$
Integrating both sides yields:
$\frac{\sigma^{2}f(t)^{2}}{2}=\tilde{C}t.$
Solving for $f(t)$, we obtain:
$f(t)=\sqrt{\frac{2\tilde{C}}{\sigma^{2}}t}.$
Let $(a_{n})$ be a sequence in $\mathbb{R}_{>0}$ defined by the recurrence relation:
$a_{n+1}=a_{n}+G(-a_{n}).$
According to Lemma 12.4, this sequence $(a_{n})$ is closely approximated by the function $f(n)$, which satisfies the associated differential equation, and we have:
$\lim_{n\to\infty}\frac{a_{n}}{f(n)}=1.$
Now, assuming that $\theta=+1$, all agents eventually choose action +1+1 starting from some time onward, with probability 1. Therefore, with probability 1, the recurrence for $\ell_{n}$ can be expressed as
$\ell_{n+1}=\ell_{n}+D_{+}(\ell_{n}),$
for all sufficiently large $n$. Additionally, by Lemma 12.2, it follows that
$\lim_{x\to\infty}\frac{D_{+}(x)}{G(-x)}=1.$
Thus, applying Lemma 12.3, we conclude that
$\lim_{n\to\infty}\frac{\ell_{n}}{a_{n}}=1,$
with probability 1. Consequently, we can combine the limits to obtain:
$\lim_{n\to\infty}\frac{\ell_{n}}{f(n)}=\lim_{n\to\infty}\frac{\ell_{n}}{a_{n}}\cdot\frac{a_{n}}{f(n)}=1,$
with probability 1, as desired. ∎
18 Proof of Theorem 4.14: Learning Rate Bounds under Smooth Randomized Response with Heterogeneous Privacy Budgets
Proof.
First, we start with the proof of the upper bound. The learning rate depends on the right-hand side of differential equation (17), i.e.,
$G_{-}(-l_{n})=\frac{1}{2}\mathbb{E}_{\varepsilon_{n}}\left[\left(e^{\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}-e^{-\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}\right)e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right].$
Define the function
$f(\varepsilon_{n})=\left(e^{\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}-e^{-\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}\right)e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}=\sinh(\varepsilon_{n})\cdot e^{\frac{\varepsilon_{n}^{2}\sigma^{2}-\varepsilon_{n}\sigma^{2}l_{n}}{2}}.$
Taking the first derivative of $f(\varepsilon_{n})$ with respect to $\varepsilon_{n}$, we apply the product rule:
$f^{\prime}(\varepsilon_{n})$ $=\frac{d}{d\varepsilon_{n}}\left[\sinh(\varepsilon_{n})\cdot e^{\frac{\varepsilon_{n}^{2}\sigma^{2}-\varepsilon_{n}\sigma^{2}l_{n}}{2}}\right]$
$=\cosh(\varepsilon_{n})\cdot e^{\frac{\varepsilon_{n}^{2}\sigma^{2}-\varepsilon_{n}\sigma^{2}l_{n}}{2}}+\sinh(\varepsilon_{n})\cdot\left(\sigma^{2}\varepsilon_{n}-\frac{\sigma^{2}l_{n}}{2}\right)\cdot e^{\frac{\varepsilon_{n}^{2}\sigma^{2}-\varepsilon_{n}\sigma^{2}l_{n}}{2}}$
$=e^{\frac{\varepsilon_{n}^{2}\sigma^{2}-\varepsilon_{n}\sigma^{2}l_{n}}{2}}\left[\cosh(\varepsilon_{n})+\left(\sigma^{2}\varepsilon_{n}-\frac{\sigma^{2}l_{n}}{2}\right)\sinh(\varepsilon_{n})\right].$
To find the optimal value $\varepsilon_{n}^{*}$ that maximizes $f(\varepsilon_{n})$, we solve
$f^{\prime}(\varepsilon_{n})=0\quad\Leftrightarrow\quad\cosh(\varepsilon_{n})+\left(\sigma^{2}\varepsilon_{n}-\frac{\sigma^{2}l_{n}}{2}\right)\sinh(\varepsilon_{n})=0.$
Since it is known that the asymptotic convergence rate is $\Theta_{\varepsilon}(\log n)$ when the privacy budget $\varepsilon$ is fixed, we now analyze whether vanishing $\varepsilon_{n}\to 0$ can lead to a faster convergence rate. To this end, we consider the behavior of $f(\varepsilon_{n})$ as $\varepsilon_{n}\to 0$, by expanding the terms in $f^{\prime}(\varepsilon_{n})$ using the Taylor expansion.
Using the Taylor series around $\varepsilon_{n}=0$, we have:
$\sinh(\varepsilon_{n})=\varepsilon_{n}+\frac{\varepsilon_{n}^{3}}{6}+o(\varepsilon_{n}^{3}),\quad\cosh(\varepsilon_{n})=1+\frac{\varepsilon_{n}^{2}}{2}+o(\varepsilon_{n}^{2}).$
Plugging these into the expression for $f^{\prime}(\varepsilon_{n})$, and keeping the terms up to first order:
$f^{\prime}(\varepsilon_{n})\approx 1-\frac{\sigma^{2}l_{n}}{2}\varepsilon_{n}=0.$
Solving this gives:
$\varepsilon_{n}^{*}\approx\frac{2}{\sigma^{2}l_{n}}\quad\text{when }l_{n}\gg 1.$
Then, by Taylor expansion, we approximate the optimal value of $f(\varepsilon_{n})$ as
$f(\varepsilon_{n}^{*})\approx\frac{2}{e\sigma^{2}l_{n}}.$
Then we need to solve the differential equation
$\frac{\mathrm{d}f}{\mathrm{d}t}(t)=G_{-}(-f(t)),$ (18)
where
$G_{-}(-f(t))=\frac{2}{e\sigma^{2}f(t)}.$
We proceed by separating variables:
$\int e\sigma^{2}f(t)\,df=2\int\,dt.$
Integrating both sides, we obtain:
$\frac{e\sigma^{2}f(t)^{2}}{2}=2t.$
Solving for $f(n)$, we find:
$f(t)=\sqrt{\frac{4t}{e\sigma^{2}}}.$
This establishes the desired upper bound $f(n)=\Theta(\sqrt{n})$.
For the lower bound, since $\mathbb{P}(\varepsilon_{n}>0)>0$, there exists a lower bound for $G_{-}(-l_{n})$
$G_{-}(-l_{n})=\frac{1}{2}\mathbb{E}_{\varepsilon_{n}}\left[\left(e^{\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}-e^{-\varepsilon_{n}+\frac{\varepsilon_{n}^{2}\sigma^{2}}{2}}\right)e^{-\varepsilon_{n}\frac{\sigma^{2}l_{n}}{2}}\right]\geq Ce^{-\tilde{\varepsilon}\frac{\sigma^{2}l_{n}}{2}}.$
where $\tilde{\varepsilon}\in(0,a]$.
Therefore, by solving the differential equation (17), we have $f(t)\geq\frac{2}{\varepsilon\sigma^{2}}\log(C\frac{\varepsilon\sigma^{2}}{2}t)$; this completes the proof.
Proof of Theorems 4.15 and 4.16: Learning Efficiency—Finite Expected Time to First Correct Action and Finite Expected Total Number of Incorrect Actions with Heterogeneous Privacy Budgets
Theorems 4.15 and 4.16 mirror the structure and argument of Theorems 4.10 and 4.11, which analyze learning efficiency in the homogeneous privacy setting. The key steps in the proof remain the same: by substituting the asymptotic learning rate result in Theorem 4.13, we obtain the corresponding expression of the expectations. Therefore, we omit the full proof and show only that the expectations in both theorems are finite.
To complete the proof, it suffices to verify that the series
$\sum_{n=1}^{\infty}e^{-\tilde{C}n^{1/2}}$
converges. Since the exponent grows sublinearly in $n$, but still diverges to infinity, this sum is dominated by an integral of the form
$\int_{1}^{\infty}e^{-\tilde{C}x^{1/2}}dx,$
which is finite by standard integral comparison. Therefore, the expected time to the first correct action $\mathbb{E}[\tau]$ and the expected total number of incorrect actions $\mathbb{E}[W]$ are both finite. ∎
19 Extension to Pufferfish Privacy
To support our findings and evaluate their robustness, we also adopt an alternative and widely studied privacy definition: Pufferfish privacy (Chatzikokolakis et al. 2013). Unlike standard differential privacy, which relies on a fixed notion of neighboring datasets, Pufferfish privacy provides a flexible framework that allows us to explicitly define which pairs of private signals should be indistinguishable. This is particularly suitable for settings with unbounded or continuous input domains, as in our case with Gaussian signals. In our context, we apply Pufferfish privacy to protect unbounded private signals by defining adjacency between signals as bounded distance. Specifically, we consider signal pairs whose distance is no greater than a predefined constant $a>0$.
Definition 19.1 ($\varepsilon$-Pufferfish privacy)
[…]
$\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=x_{n})\leq\exp(\varepsilon)\,\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=x_{n}).$ (19)
To satisfy the above privacy guarantee and maximize the expected utility, we show a staircase variant of the randomized response mechanism will achieve them, where the flipping probability is softly adapted to the signal value.
Definition 19.2 (Staircase Randomized Response)
A strategy $\mathcal{M}(s_{n};h_{n-1})$ is called a staircase randomized response strategy with flip probability $u_{n}$ if, given an initial decision $a_{n}$, the reported action $x_{n}$ is determined as follows:
$\mathbb{P}_{\mathcal{M}}(x_{n}=+1\mid a_{n}=-1)=\mathbb{P}_{\mathcal{M}}(x_{n}=-1\mid a_{n}=+1)=u_{n}(s_{n}),$
$\mathbb{P}_{\mathcal{M}}(x_{n}=+1\mid a_{n}=+1)=\mathbb{P}_{\mathcal{M}}(x_{n}=-1\mid a_{n}=-1)=1-u_{n}(s_{n}),$
with the flip probability $u_{n}$ defined as:
$u_{n}(s_{n})=\frac{e^{-k\varepsilon}}{1+e^{\varepsilon}},\quad\text{if }s_{n}\in(t(l_{n})+ka,\ t(l_{n})+(k+1)a)\cup(t(l_{n})-(k+1)a,\ t(l_{n})-ka),\ k\in\mathbb{N}.$
Here, $t(l_{n})$ represents the threshold value for different actions as a function of the public belief $l_{n}$. If $s_{n}>t(l_{n})$, agents prefer to choose action +1+1 before flipping the action. Conversely, if $s_{n}<t(l_{n})$, agents prefer to choose action 1-1 before flipping. The mechanism is symmetric around $t(l_{n})$, and the flip probability decays exponentially with the distance from the threshold.
Theorem 19.3
In the sequential learning model with Gaussian signals, the optimal reporting strategy under a fixed privacy budget $\varepsilon$ is the staircase randomized response mechanism. This strategy satisfies $\varepsilon,$-pufferfish privacy.
Proof.
To derive an upper bound for the public log-likelihood ratio under the staircase randomized response mechanism, we analyze the evolution of the public belief through the log-likelihood ratio. In the absence of privacy constraints, agent $n$ chooses $a_{n}=+1$ if and only if
$l_{n}+L_{n}>0.$
This implies that there exists a threshold $t(l_{n})$, determined by the public belief $l_{n}$, such that the agent selects action $a_{n}=+1$ if $s_{n}>t(l_{n})$, and $a_{n}=-1$ otherwise.
We first consider the case where both signals $s_{n},s_{n}^{\prime}\in(t(l_{n})-a,\ t(l_{n})+a)$. Since the flipping probability in this region is set to $u_{n}(s_{n})=\frac{1}{1+e^{\varepsilon}}$, the privacy loss is bounded as
$\frac{\mathbb{P}(\mathcal{M}(s_{n};h_{n-1})=-1)}{\mathbb{P}(\mathcal{M}(s_{n}^{\prime};h_{n-1})=-1)}\leq e^{\varepsilon}.$
Next, consider the general case where $s_{n},s_{n}^{\prime}\in(t(l_{n})+(k-1)a,\ t(l_{n})+(k+1)a)\cup(t(l_{n})-(k+1)a,\ t(l_{n})-(k-1)a)$ for $k\in\mathbb{N}$. According to the definition of the staircase randomized response, the flipping probability satisfies
$u_{n}(s_{n})=\frac{e^{-k\varepsilon}}{1+e^{\varepsilon}}.$
Thus, the privacy ratio between adjacent signals in this interval is bounded by $e^{\varepsilon}$, ensuring that the mechanism satisfies $\varepsilon$-Pufferfish privacy.
To show the optimality of this mechanism, we note that for $s_{n},s_{n}^{\prime}\in(t(l_{n})-a,\ t(l_{n})+a)$, the two signals lead to distinct actions. Therefore, to satisfy $\varepsilon$-Pufferfish privacy, the minimal flipping probability must be at least $\frac{1}{1+e^{\varepsilon}}$. Similarly, for $s_{n},s_{n}^{\prime}\in(t(l_{n})+(k-1)a,\ t(l_{n})+(k+1)a)$ or its symmetric counterpart on the left, the minimal flipping probability required is
$u_{n}(s_{n})=\frac{e^{-k\varepsilon}}{1+e^{\varepsilon}}.$
If one chooses any flipping probability smaller than this, the privacy guarantee would be violated. Therefore, the staircase randomized response is the optimal strategy under this privacy constraint. ∎
Compared to the smooth randomized response defined under metric differential privacy, the staircase randomized response exhibits the same asymptotic behavior in terms of flip probability: both decay exponentially as the signal deviates from the decision threshold. For the staircase randomized response, the flip probability $u_{n}(s_{n})$ satisfies the following bounds:
$\frac{1}{1+e^{\varepsilon}}\,e^{-\frac{\varepsilon}{a}|s_{n}-t(l_{n})|}\leq u_{n}(s_{n})\leq\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\,e^{-\frac{\varepsilon}{a}|s_{n}-t(l_{n})|}.$
As a result, all theoretical guarantees established for the continuous signal model under the smooth mechanism continue to hold under the staircase mechanism, as long as the step size $a$ is a finite constant. Since the structure of the proofs remains largely unchanged, we omit detailed derivations and directly present the theorems under the Pufferfish privacy framework.
Theorem 19.4 (Learning Rate under Staircase Randomized Response )
Consider a fixed differential privacy budget $\varepsilon$ and a smooth randomized response strategy for sequential learning with Gaussian signals. Define the constant $C_{\varepsilon,a}$ as bounded by
$\frac{\varepsilon\sigma^{2}}{2a(1+e^{\varepsilon})}\left(e^{\frac{\varepsilon}{a}+\frac{\varepsilon^{2}\sigma^{2}}{2a^{2}}}-e^{-\frac{\varepsilon}{a}+\frac{\varepsilon^{2}\sigma^{2}}{2a^{2}}}\right)\leq C_{\varepsilon,a}\leq\frac{\varepsilon e^{\varepsilon}\sigma^{2}}{2a(1+e^{\varepsilon})}\left(e^{\frac{\varepsilon}{a}+\frac{\varepsilon^{2}\sigma^{2}}{2a^{2}}}-e^{-\frac{\varepsilon}{a}+\frac{\varepsilon^{2}\sigma^{2}}{2a^{2}}}\right).$
Then, for any $\frac{\varepsilon}{a}\in(0,\infty)$, the convergence rate of the public log-likelihood ratio satisfies
$f(n)=\frac{2a}{\varepsilon\sigma^{2}}\log(C_{\varepsilon,a}\cdot n)\sim\frac{2a}{\varepsilon\sigma^{2}}\log(n),\quad\text{as }n\to\infty.$
When the parameter $a$ increases, if the ratio $\frac{a}{\varepsilon}$ is finite, then the asymptotic learning rate $f(n)=\frac{2a}{\varepsilon\sigma^{2}}\log(C_{\varepsilon,a}n)$ increases. This is because a larger $a$ induces a more asymmetric smooth randomized response, which improves the informativeness of the observed actions and thus accelerates the overall learning process. However, this improvement comes with a caveat: as $a\to\infty$, the term $C_{\varepsilon,a}$ tends to zero, which causes $f(n)$ to diverge. Therefore, we restrict our attention to the regime where $\frac{\varepsilon}{a}$ is constant and bounded away from zero to ensure that the learning rate remains well defined and informative.
Theorem 19.5 (Finite Expected Time to First Correct Action under Pufferfish Privacy)
Consider a fixed differential privacy budget and suppose that agents follow a smooth randomized response strategy under Gaussian signals in a sequential learning setting. Then, the expected time $\mathbb{E}[\tau]$ until the first correct action satisfies
$\mathbb{E}[\tau]=C_{1}C_{\varepsilon,a}^{-\frac{2a}{\varepsilon\sigma^{2}}}\zeta\left(\frac{2a}{\varepsilon\sigma^{2}}\right),$
where $C_{1}$ is a positive constant that does not depend on $\varepsilon$ and $\zeta\left(x\right)=\sum_{n=1}^{\infty}n^{-x}$. If $\varepsilon<\frac{2a}{\sigma^{2}}$, the series converges and thus the expected time to the first correct action is finite:
$\mathbb{E}[\tau]<+\infty.$
Theorem 19.6 (Finite Expected Total Number of Incorrect Actions under Pufferfish Privacy)
Consider a fixed differential privacy budget and suppose that agents follow a smooth randomized response strategy under Gaussian signals in a sequential learning setting. Then, the expected total number of incorrect actions $\mathbb{E}[W]$ satisfies
$\mathbb{E}[W]=C_{2}C_{\varepsilon,a}^{-\frac{2a}{\varepsilon\sigma^{2}}}\zeta\left(\frac{2a}{\varepsilon\sigma^{2}}\right),$
where $C_{2}$ is a positive constant that does not depend on $\varepsilon$. If and only if $\varepsilon<\frac{2a}{\sigma^{2}}$, the series converges and thus the expected total number of incorrect actions is finite:
$\mathbb{E}[W]<+\infty.$
Although a higher value of $a$ makes it easier to ensure the finiteness of both $\mathbb{E}[\tau]$ and $\mathbb{E}[W]$, this benefit is not without potential drawbacks. In particular, as $a\to\infty$, the term $C_{\varepsilon,a}^{-\frac{2a}{\varepsilon\sigma^{2}}}$ diverges to infinity. Therefore, our results are valid only under the condition that $\frac{\varepsilon}{a}$ is constant and bounded away from zero. This ensures that the learning efficiency bounds remain meaningful and finite.